### v2

# Concepts in Motion: Temporal Bottlenecks for Interpretable Video Classification

Abstract
Concept Bottleneck Models (CBMs) enable interpretable image classification by structuring predictions around human-understandable concepts, but extending this paradigm to video remains challenging due to the difficulty of extracting concepts and modeling them over time. In this paper, we introduce MoTIF (Moving Temporal Interpretable Framework), a transformer-based concept architecture that operates on sequences of temporally grounded concept activations, by employing per-concept temporal self-attention to model when individual concepts recur and how their temporal patterns contribute to predictions. Central to the framework is an agentic concept discovery module to automatically extract object- and action-centric textual concepts from videos, yielding temporally expressive concept sets without manual supervision. Across multiple video benchmarks, this combination substantially narrows the performance gap between interpretable and black-box video models while maintaining faithful and temporally grounded concept explanations. Code available at github.com/patrick-knab/MoTIF.
Machine Learning, ICML
1 Introduction
Modern deep learning models achieve outstanding results in video tasks such as video classification, action recognition, and event detection (Liu et al., 2021; Bertasius et al., 2021). Despite their success, these models are commonly perceived as black boxes since their internal workings are not interpretable (Molnar et al., 2020; Knab et al., 2025b) and this opacity can limit deployment in medical domains, where interpretability may matter more than peak accuracy. Concept Bottleneck Models (CBMs) (Koh et al., 2020) address this issue by enforcing an intermediate bottleneck layer of human-understandable concepts, which are then used by a linear classifier to generate the final prediction.
Figure 1: Overview of MoTIF. Video classification pipeline with agentic concept discovery and its three explanation modes.
Figure 2: MoTIF. The framework takes videos as input and produces local concept explanations for local windows, global explanations for entire videos, and temporal dependency maps from the attention heads of the transformer module. Model represents MoTIF (ViT-L14) and sample frames are from HMDB51 (Kuehne et al., 2011), licensed under $C$ BY 4.0.
While CBMs have been extensively studied in the image domain (Prasse et al., 2025; Yang et al., 2023; Sun et al., 2025; Schrodi et al., 2025), their extension to video remains largely unexplored (Jeyakumar et al., 2022). Videos differ from images in that they contain a temporal component: concepts evolve over time, and many actions cannot be inferred from a single frame (Lee et al., 2025b; Chen et al., 2025). Importantly, the challenge of variable video length is orthogonal to modeling long-range dependencies. While transformers (Vaswani et al., 2017) excel at capturing such dependencies (Bertasius et al., 2021), their dense temporal feature mixing obscures concept-level attributions, limiting their suitability for interpretable concept bottleneck modeling (Hao et al., 2021; Molnar et al., 2020). This highlights the need for architectures that handle variable-length inputs while preserving concept-level interpretability over time.
In this work, we introduce MoTIF (Moving Temporal Interpretable Framework), a concept bottleneck model tailored for video classification. The CBM is not designed to replace high-capacity black-box video models, but to make temporal concept reasoning explicit and inspectable. MoTIF builds on transformer-inspired blocks and introduces a per-channel temporal self-attention (diagonal attention) mechanism that isolates temporal reasoning for each concept to enable interpretation. In addition, we utilize an agentic concept discovery approach to generate object- and action-based textual, domain-specific concepts (Fig. 1). This setup removes the need for manually curated concept vocabularies, which is particularly valuable in video domains where relevant objects and actions vary substantially across datasets. To illustrate how MoTIF extends beyond static images, Fig. 2 shows our framework: it processes a video, tracks its concepts through time, and explains which concepts drive the final prediction. In this example, the model identifies recurring semantic motifs such as a bow being mounted and the arrow being shot by Legolas (Lord of the Rings), which together form the basis for recognizing the higher-level action “shooting a bow.” Temporal dependency maps further reveal that specific concepts primarily attend to characteristic frames, i.e., the moments when the bow is mounted and the arrow released. As this example shows, the design preserves concept separation across time and enables analysis of both global concept contributions over an entire video and local relevance within specific frames.
Our key contributions are:
  • MoTIF, a Video CBM with per-channel temporal self-attention that preserves concept independence within Transformer blocks while modeling temporal dynamics at the level of individual concepts.
  • An unsupervised agentic concept discovery pipeline that extracts object and action concepts directly from video and substantially improves the performance of interpretable video CBMs, narrowing the gap to black-box models.
  • MoTIF enables three complementary explanation modes: (i) global concept relevance via log-sum-exp (LSE) pooling, (ii) localized temporal explanations using windowed concept attributions, and (iii) attention-based temporal maps over concept activations.
2 Method
In this section, we present the agentic concept discovery and the MoTIF framework for interpretable video classification. The goal is to preserve concept-level transparency while leveraging temporal modeling capabilities.
MoTIF is a CBM with a transformer architecture for video classification that operates directly on sequences of per-window concept activations (Figure 2). A video is represented as a sequence of $T$ temporal windows, each encoded by $C$ scalar concept activations (one per concept), where $C=|\mathcal{C}|$ is given by the concept bank produced by the agentic discovery module.
The framework enforces strict and intentional per-concept interpretability: temporal dependencies are modeled independently for each concept channel, recurrent temporal patterns—motifs—remain attributable to individual concepts. After temporal processing, per-concept activations are refined via a nonnegative affine transformation, and a classifier produces per-time-step logits. Video-level predictions are obtained via log-sum-exp pooling, which also yields a time-importance profile for additional explanation. In contrast, a conventional transformer mixes channels during attention (Vaswani et al., 2017)  [^1] We also evaluate full multi-head attention, but emphasize that this removes explicit concept attribution. A detailed analysis and other design choices are provided in Section 3.2.2 and Appendix C. For completeness, Appendix D provides an algorithmic overview of MoTIF, including its main components and the concept discovery module.
Channels as concepts. We use the term channel to denote one dimension of the concept space. Each video $n$ yields $y^{(n)}\in\mathbb{R}$ classes, $X^{(n)}\in\mathbb{R}^{T_{n}\times C}$ concept activations, where the second axis ($C$) indexes concepts; each slice $X^{(n)}_{:,c}\in\mathbb{R}^{T_{n}}$ corresponds to the temporal activation sequence of concept $c\in C$. Unlike CNN channels, these dimensions are semantically interpretable by design.
Notation. Capital and lowercase letters denote matrices and vectors, respectively. During training, a batch of videos is represented as $X\in\mathbb{R}^{B\times T\times C}$, where $B$ is the batch size, $T$ is the (padded) sequence length, and $C$ is the number of concepts. Thus $X_{:,:,c}\in\mathbb{R}^{B\times T}$ contains the activations of concept $c$ across the batch.
2.1 Agentic Concept Discovery
The agentic concept discovery stage converts raw visual inputs into structured, interpretable concept activations that serve as input to the MoTIF backbone. While formulated for video, images are naturally handled as the special case of a single temporal window.
Temporal windowing. Given a video $X=\{I_{t}\}_{t=1}^{T}$, we partition it into $n$ temporal windows
$\mathcal{W}(x)=\{w_{k}\}_{k=1}^{n},\qquad w_{k}=\{I_{t}\}_{t=s_{k}}^{s_{k}+W-1},$ (1)
where $W$ denotes the window length and $\{s_{k}\}$ are the start indices. The number of windows $K$ is determined by the video duration, window size, and overlap. Smaller windows favor short-lived, fine-grained events, while larger windows emphasize persistent semantic content. This design provides explicit control over the temporal resolution at which concepts are discovered.
Agentic concept proposal. Each window $w_{k}$ is processed independently by a vision–language agent $A_{\mathrm{vlm}}$, which generates natural-language concept candidates
$(\mathcal{C}^{\text{obj}}_{w_{k}},\mathcal{C}^{\text{act}}_{w_{k}})=A_{\mathrm{vlm}}(w_{k}),$ (2)
capturing objects, attributes, actions, and interactions. This step is task-agnostic and does not rely on any predefined concept vocabulary (Appendix E.1).
Concept and video embeddings. The global concept bank $\mathcal{C}$ is constructed from the union of all discovered textual object and action concepts. To ensure semantic diversity and remove near-duplicates, concepts are embedded using the same image–text aligned backbone as MoTIF (CLIP), and filtered by cosine similarity: concepts with pairwise similarity greater than $0.9$ are removed. For each window, depending on the backbone, visual embeddings are obtained either from a representative frame or a video-adapted CLIP variant (Appendix A.3). Concept activations are computed as cosine similarities between window embeddings and concept embeddings, yielding per-video activation matrices
$X^{(n)}\in\mathbb{R}^{T_{n}\times C},$ (3)
where each channel corresponds to one interpretable concept and is processed independently by MoTIF.
2.2 Transformer Bottleneck Model
2.2.1 Per-Channel Temporal Self-Attention
In standard transformers, query–key–value (QKV) projections are implemented as full linear layers ($W_{Q}\in\mathbb{R}^{C\times C}$), which mix channels and would obscure concept attribution. In MoTIF, we avoid mixing: each concept $c$ receives its own QKV projections via depthwise $1{\times}1$ convolutions,
$Q,K,V$ $\in\mathbb{R}^{T\times C},Q_{:,c}$ $=x_{:,c}\,\theta_{Q}^{(c)},$ (4)
$K_{:,c}$ $=x_{:,c}\,\theta_{K}^{(c)},V_{:,c}$ $=x_{:,c}\,\theta_{V}^{(c)}.$
Here $\theta_{Q}^{(c)},\theta_{K}^{(c)},\theta_{V}^{(c)}\in\mathbb{R}$ are channel-specific scalars applied uniformly across all $T$, not temporal filters. Equivalently, each projection uses a depthwise kernel $\Theta_{Q},\Theta_{K},\Theta_{V}\in\mathbb{R}^{C\times 1\times 1}$, where $\theta_{*}^{(c)}$ is the $c$-th depthwise filter. Thus, temporal information is preserved and channels remain independent, without cross-channel mixing.
Attention scores are computed per concept as $W_{c,t,u}\;=\;Q_{t,c}\,K_{u,c}$, where $t$ denotes the query time step and $u$ the key/value time step. A softmax over $u$ yields attention weights $W\in\mathbb{R}^{C\times T\times T}$, so that each concept decides which of its past or future activations to attend to. The output is obtained as the weighted sum over $V$:
$X^{(L)}_{t,c}=\sum_{u=1}^{T}W_{c,t,u}\,V_{u,c}.$ (5)
The difference between full and diagonal attention is visualized in Equation 6 as the structure of the concept–concept attention matrix. This is why we call this mechanism diagonal attention: each channel learns to construct its own temporal filter by weighting past activations differently at each step. Unlike fixed convolutional kernels, the attention weights adapt to the input sequence, yet the restriction to depthwise (per-concept) projections guarantees that evidence for one concept does not leak into another.
$\begin{array}[]{cc}\underbrace{\left[\begin{smallmatrix}(c_{1}\to\ c_{1})&\cdots&(c_{1}\to\ c_{C})\\ \vdots&\ddots&\vdots\\ (c_{C}\to\ c_{1})&\cdots&(c_{C}\to\ c_{C})\end{smallmatrix}\right]}_{\text{Full attention}}&\underbrace{\left[\begin{smallmatrix}(c_{1}\to\ c_{1})&&\\ &\ddots&\\ &&(c_{C}\to\ c_{C})\end{smallmatrix}\right]}_{\text{Diagonal attention}}\end{array}$ (6)
The block concludes with per-channel normalization and a lightweight feed-forward network (two depthwise $1{\times}1$ convolutions with GELU and dropout). In this work, we deliberately isolate temporal reasoning per concept to preserve attribution. Importantly, this is a design choice rather than a limitation of the framework: as shown in Section 3.2.2 enabling cross-concept interactions improves accuracy at the cost of interpretability, highlighting a controllable trade-off rather than a hard restriction.
Architectural extension. For most experiments in this paper, we report the results of MoTIF using diagonal attention within a standard transformer architecture. However, as shown by (Bertasius et al., 2021), separating spatial and temporal attention can further enhance performance. Therefore, we additionally evaluate MoTIF when extended to a space-time transformer architecture (MoTIF-ST), as detailed in Appendix A.4, demonstrating that MoTIF is not restricted to a single transformer design.
Complexity. Diagonal attention reduces the channel-mixing cost from $\mathcal{O}(C^{2}T)$ to $\mathcal{O}(CT)$, but it requires computing a full $T\times T$ attention map for every channel, yielding $\mathcal{O}(CT^{2})$. By contrast, standard multi-head attention only maintains a fixed number of maps (one per head), scaling as $\mathcal{O}(HT^{2})$ with $H\ll C$. Thus, with many concepts, full attention is often more efficient, while diagonal attention deliberately trades computational cost for strict concept isolation.
2.2.2 Per-Concept Affine Transformation
Each refined activation $X^{(L)}_{t,c}$ can optionally be scaled and shifted by learnable concept-specific parameters, $\tilde{X}_{t,c}=\gamma_{c}X^{(L)}_{t,c}+\delta_{c}$, and then passed through a Softplus nonlinearity $Z_{t,c}=\mathrm{Softplus}(\tilde{X}_{t,c})$, which ensures nonnegative concept activations while avoiding dead units and maintaining differentiability everywhere. This transformation introduces a per-concept scale ($\gamma_{c}$) and bias ($\delta_{c}$), allowing the model to adapt to differences in concept magnitude and activation thresholds.
2.2.3 Classification Head
From these activations, per-time-step logits are computed as $\ell_{t}=W_{k}Z_{t,:}+b$ with $W_{k}\in\mathbb{R}^{K\times C}$, where $K$ denotes the number of target classes and $C$ the number of concepts. Since videos vary in length, we apply log-sum-exp (LSE) pooling across time (Wang et al., 2018), which smoothly interpolates between mean-pooling($\tau\to\ 0$) and max-pooling ($\tau\to\ \infty$):
$\hat{c}=\tfrac{1}{\tau}\log\sum_{t=1}^{T}m_{t}e^{\tau c_{t}},\qquad\hat{\ell}=\tfrac{1}{\tau}\log\sum_{t=1}^{T}m_{t}e^{\tau\ell_{t}},$ (7)
where $m_{t}\in{0,1}$ are masks for padded windows. We denote the pooled concept vector by $\hat{c}$ and the pooled logits by $\hat{\ell}$. The pooled logits $\hat{\ell}$ form the video-level prediction.
Training objective. The model is trained with class-weighted cross-entropy on $\hat{\ell}$, complemented with two regularizers: an $\ell_{1}$ penalty on $W$ to encourage sparsity, and an activation sparsity penalty on $Z$:
$\mathcal{L}$ $=\mathrm{CE}(\hat{\ell},y)+\lambda_{\ell_{1}}\,\lVert W_{k}\rVert_{1}$ (8)
$\quad+\lambda_{\mathrm{sparse}}\;\frac{1}{(\sum_{t}m_{t})C}\sum_{t,c}m_{t}\lvert Z_{t,c}\rvert$.
2.3 Explanation
To make predictions transparent, MoTIF decomposes them into time- and concept-resolved contributions. For a target class $k$, the contribution of each time step is $c_{t}^{(k)}=Z_{t,:}W_{k,:}$, with score $s_{t}^{(k)}=\sum_{c=1}^{C}c_{t,c}^{(k)}+b_{k}$. Temporal importance weights are derived consistently with LSE pooling:
$\pi_{t}^{(k)}=\frac{\exp(s_{t}^{(k)}/\tau)}{\sum_{u=1}^{T}\exp(s_{u}^{(k)}/\tau)}.$ (9)
Aggregating over time yields global concept attributions, $\bar{c}^{(k)}=\sum_{t=1}^{T}\pi_{t}^{(k)}\,c_{t}^{(k)}$. MoTIF therefore provides three complementary views: (1) Global concepts, via $\bar{c}^{(k)}$; (2) Local concepts, active in high-weight windows (large $\pi_{t}^{(k)}$); (3) Temporal dependencies, revealed by per-concept attention maps ($W_{c,t,u}$) that show how occurrences of concepts relate across time (see Appendix B.3). Together, these expose both which concepts mattered and when they were decisive.
Table 1: Performance comparison (% Top-1 accuracy). Mean $\pm$ standard deviation on train-test splits across datasets, using agentic concept discovery for MoTIF, and comparing against zero-shot, global CBM, and black-box baselines.
Method Breakfast HMDB51 UCF101 SSv2
Zero-shot CLIP-RN/50 (Radford et al., 2021) 18.6 $\pm$ 2.6 29.8 $\pm$ 0.5 57.2 $\pm$ 0.9 0.8
CLIP-ViT-B/32 (Radford et al., 2021) 23.2 $\pm$ 2.9 38.1 $\pm$ 0.3 59.9 $\pm$ 0.4 0.9
CLIP-ViT-L/14 (Radford et al., 2021) 31.1 $\pm$ 4.7 45.7 $\pm$ 0.1 70.6 $\pm$ 0.5 0.9
PE-L/14 (Bolya et al., 2025) 41.4 $\pm$ 7.0 56.7 $\pm$ 0.6 74.6 $\pm$ 0.9 2.2
PE-G/14 (Bolya et al., 2025) 47.4 $\pm$ 5.4 60.7 $\pm$ 1.0 74.6 $\pm$ 0.9 2.2
Global CBM CLIP-RN/50 (Radford et al., 2021) 36.9 $\pm$ 7.7 61.8 $\pm$ 1.7 84.1 $\pm$ 0.8 17.0
CLIP-ViT-B/32 (Radford et al., 2021) 38.3 $\pm$ 9.5 62.6 $\pm$ 0.8 86.4 $\pm$ 0.6 18.2
CLIP-ViT-L/14 (Radford et al., 2021) 57.0 $\pm$ 8.0 71.0 $\pm$ 1.1 93.4 $\pm$ 0.7 22.0
PE-L/14 (Bolya et al., 2025) 72.4 $\pm$ 8.3 76.4 $\pm$ 0.8 96.3 $\pm$ 0.1 31.3
PE-G/14 (Bolya et al., 2025) 75.8 $\pm$ 7.1 77.8 $\pm$ 0.8 97.5 $\pm$ 0.4 33.6
MoTIF (ours) MoTIF (RN/50) 55.1 $\pm$ 7.1 66.2 $\pm$ 0.5 86.7 $\pm$ 0.6 20.0
MoTIF (ViT-B/32) 52.7 $\pm$ 5.8 68.5 $\pm$ 1.0 88.5 $\pm$ 0.6 20.7
MoTIF (ViT-L/14) 71.0 $\pm$ 6.2 76.1 $\pm$ 0.5 94.8 $\pm$ 0.5 25.8
MoTIF-ST (ViT-L/14) 72.6 $\pm$ 6.5 75.8 $\pm$ 0.6 94.8 $\pm$ 0.4 27.7
MoTIF (PE-L/14) 83.2 $\pm$ 6.2 81.8 $\pm$ 0.6 97.0 $\pm$ 0.3 37.3
MoTIF-ST (PE-L/14) 85.4 $\pm$ 6.3 80.8 $\pm$ 1.0 97.2 $\pm$ 0.2 39.6
MoTIF (PE-G/14) 87.5 $\pm$ 4.9 83.0 $\pm$ 0.6 98.0 $\pm$ 0.2 40.4
MoTIF-ST (PE-G/14) 87.3 $\pm$ 7.1 82.1 $\pm$ 1.0 98.4 $\pm$ 0.3 41.9
Black- Box TSM (Lin et al., 2019) 59.1 73.5 95.9 61.7
No frame left behind (Liu et al., 2021) 62.0 73.4 96.4 62.7
VideoMAE V2 (Wang et al., 2023) 88.1 99.6 76.8
3 Experiments
In this section, we first evaluate MoTIF on common video classification benchmarks. We then analyze different architectural design choices through ablation studies, illustrate the kinds of explanations our model provides with representative examples, and finally examine its behavior under interventions.
3.1 Experimental Setup
Datasets. We evaluate MoTIF on four established benchmarks: Breakfast Actions (Kuehne et al., 2014), HMDB51 (Kuehne et al., 2011), UCF101 (Soomro et al., 2012), and Something-Something V2 (SSv2) (Goyal et al., 2017; Materzynska et al., 2020). These datasets contain 10, 51, 101, and 174 classes, respectively. These datasets span diverse settings: short versus long clips, local versus global temporal dependencies, and varying degrees of abstraction and viewpoint diversity.
Backbones. We adopt a range of CLIP-based backbones differing in model size, patch resolution, and temporal adaptation as feature extractors. Specifically, we include RN/50, ViT-B/32, and ViT-L/14 from CLIP (Radford et al., 2021), SigLIP (Zhai et al., 2023), and video-adapted Perception Encoder (PE) (Bolya et al., 2025). SigLIP replaces CLIP’s contrastive softmax with a pairwise sigmoid loss, decoupling image and text encoders, which improves training efficiency and scalability. The Perception Encoder is trained on video–text pairs and incorporates a mechanism to aggregate frame-level information. For SigLIP, we use ViT-L/14; for PE, we use both ViT-L/14 and ViT-G/14.
Concept Discovery. For agentic concept discovery (Section 2.1), we use Qwen-3 30B (Yang et al., 2025) to generate the concept sets employed in our main experiments in Table 1. For each dataset, up to five training videos per class are provided to the agent, and concept extraction is performed exclusively on the training split to avoid information leakage. For ablation studies, we follow Yang et al. (2023) and construct smaller, dataset-specific concept sets by prompting an LLM (GPT-5) (OpenAI et al., 2024). Details on the resulting concept vocabularies are provided in Appendix E.
Baselines. In line with (Rao et al., 2024; Prasse et al., 2025), we evaluate MoTIF against zero-shot and linear-probe baselines. To ensure a fair comparison with MoTIF, we compute zero-shot predictions on the window-level and apply majority voting across windows, preventing the zero-shot baseline from relying solely on a single global aggregation. As a supervised baseline (Global CBM), we train a linear classifier on top of mean-pooled window embeddings, analogous to a global bottleneck. This setup captures only coarse video-level information, in contrast to MoTIF, which models temporally localized concept activations. All experiments were conducted on a single H100 GPU with 94 GB of VRAM.
3.2 Experimental Results
3.2.1 Performance Evaluation
We report Top-1 accuracies with different backbones across four datasets in Table 1, using the tuned hyperparameters described in Appendix A.1 and agentic concept discovery. The global CBM consistently outperforms zero-shot, while MoTIF surpasses both. Accuracy increases with backbone capacity, suggesting that stronger vision–language representations yield more informative (and thus more easily separable) concept activations, which improves the downstream linear classifier. This indicates that introducing a concept bottleneck can preserve—and sometimes improve—predictive performance while adding interpretability.
Regarding the transformer design, both variants perform well; however, the space–time variant (MoTIF-ST) (Bertasius et al., 2021) consistently achieves higher accuracy on SSv2. Across comparable model sizes, PE-based backbones outperform CLIP variants. Overall performance is strong across datasets: On Breakfast, MoTIF improves over the global CBM by 11111818 percentage points. For HMDB51, UCF101, and SSv2, gains are smaller (0.90.98.28.2 points), likely because these datasets contain short clips where mean pooling is often sufficient—particularly for PE embeddings, which yield the strongest video representations. A performance gap between interpretable concept bottleneck models and strong non-interpretable black-box baselines is primarily observed for VideoMAE V2 (Wang et al., 2023); for TSM (Lin et al., 2019) and NoFrameLeftBehind (Liu et al., 2021), results are similar or better except on SSv2. SSv2 isolates fine-grained relational reasoning between multiple concepts, a setting where enforcing strict concept independence intentionally trades expressivity for interpretability. Despite its explicit focus on interpretability, MoTIF substantially narrows the accuracy gap to black-box models. This is achieved by combining temporally grounded concept modeling with agentic concept discovery, which yields more complete and task-relevant concept sets; an overview of the LLM-based concept set (cf. Table 9) and the corresponding test-split results are provided in Appendix C.1.
3.2.2 Ablations
We perform a series of ablation studies to evaluate the effect of key design choices in MoTIF. Unless stated otherwise, experiments on Breakfast Actions and HMDB51 use CLIP ViT-B/32 and RN/50 backbones (hyperparameters in Appendix A.1). These two were chosen to cover distinct inductive biases: ViT-B/32 as a transformer-based (Dosovitskiy et al., 2020) extractor and RN/50 as a ResNet CNN (He et al., 2016), allowing evaluation of MoTIF across architecture families. These two datasets differ in scale, variability, and action granularity, making them complementary testbeds for analyzing architectural contributions. Further ablations can be found in Appendix C.
Figure 3: Effect of log-sum-exp temperature $\tau$ on accuracy and entropy. Accuracy remains stable, while concept- and logit-level entropy decrease as $\tau$ increases.
Temperature $\tau$. We vary the log-sum-exp pooling temperature $\tau$ to assess its effect on both accuracy and the sharpness of the time-importance distribution. Sharpness is quantified by the entropy of the softmax weights, computed either at the concept or at the logits level. Experiments show (see Figure 3) that accuracy varies only slightly across all tested values of $\tau$, indicating robustness of the predictive performance. In contrast, entropy decreases monotonically with larger $\tau$, i.e., the temporal weighting becomes more concentrated on fewer decisive frames. Hence, $\tau$ provides a controllable parameter: small values result in diffuse, high-entropy explanations, whereas large values produce sharp, low-entropy attributions, while accuracy remains relatively stable. Thus this parameter can be tuned by the user.
Figure 4: Full vs. diagonal attention. Train and test accuracy with and without enforcing diagonal attention over five seeds.
Attention variant (full vs. diagonal). While MoTIF enforces concept isolation, we also evaluate a variant with full multi-head attention. Figure 4 compares train and test accuracy for both settings. This comparison explicitly evaluates the impact of cross-concept interactions on both predictive performance and interpretability. Diagonal attention exposes a clear accuracy–interpretability trade-off: enabling cross-concept attention recovers accuracy on demanding datasets, while diagonal attention preserves faithful concept attribution, allowing controlled analysis of how temporal interactions affect both prediction and explanation. On temporally demanding datasets such as SSv2, this trade-off becomes more pronounced, with full attention yielding up to 10.1% higher test accuracy. We further illustrate the effect on explanations in Figure B.2: when channels are mixed, the most influential dimensions appear effectively random, substantially reducing interpretability.
Temporal sensitivity and dynamic concepts. A general challenge in video concept learning is preventing models from defaulting to temporally order-invariant reasoning (Bertasius et al., 2021). To explicitly assess whether MoTIF captures temporal structure, we evaluate temporal dependence in two complementary ways. First, we construct a controlled synthetic benchmark (1,989 sequences, matching the number of Breakfast training instances) in which class identity depends exclusively on frame ordering (ascending, descending, U-shaped, inverted-U, and periodic patterns), such that static appearance cues are uninformative by design (for details see Appendix B.4). Performance collapses from 86.97% on ordered sequences to 21.06% under temporal shuffling (chance $\approx 20\%$), demonstrating that MoTIF explicitly encodes temporal order rather than relying on static appearance. Moreover, on this benchmark the static image-level bottleneck baseline (Global CBM), achieves only 35.5% accuracy, underscoring that temporal modeling is essential when order defines the class.
Table 2: Temporal sensitivity analysis and bottleneck comparison (PE-L/14). Shuffling indicates random permutation of windows at evaluation time; synthetic uses five temporal classes.
Setting Basic Shuffled
Synthetic (MoTIF) 86.97 21.06
Synthetic (Global CBM) 35.5 35.5
Breakfast (MoTIF) 87.3 85.2–86.6
HMDB51 (MoTIF) 79.9 77.1–78.5
UCF101 (MoTIF) 94.7 94.5–94.6
SSv2 (MoTIF) 30.0 26.9–27.4
Second, in Table 2, we apply the same shuffling intervention to real-world datasets. Overall, the degradation is mild across all benchmarks (typically a few percentage points), suggesting that many clips still contain strong appearance cues. SSv2 shows the largest drop (about 2–3 points), which is consistent with its higher reliance on temporal relations; correspondingly, it benefits the most from agent-generated temporal concepts.
3.2.3 Explanations
Figure 5: MoTIF explanations. Example videos from Breakfast and UCF101 with correct classifications, illustrating the three explanation modes supported by MoTIF (ViT-L14).
A central goal of our approach is to provide interpretable insights into the decision-making process of video classification models. We present further examples for the three explanation modes introduced in Section 2.3: global concept importance, local concepts in decisive windows, and temporal dependencies between concepts. In Figure 5, the first example from Breakfast Action shows preparing a sandwich. Global and temporal concepts highlight bread and bagel. The attention map for bagel reveals strong vertical stripes, meaning many query frames attend to the same key frames—anchor moments where the concept is most pronounced. Thus, bagel is localized in time but influences the entire sequence. The second example, from UCF101, depicts kayaking. Here, the most salient concepts are paddle and kayak, which remain stable over the short clip. The attention map for paddle shows a more uniform distribution across time, with slight emphasis at $u=3$ and $u=6$, indicating consistent expression of the concept rather than isolated peaks. Additional video examples are provided in the supplementary material (Appendix B.1 discusses misclassifications, and B.2 shows examples without diagonal attention); all data will be released upon acceptance.
3.2.4 Concept Interventions
Table 3: Concept interventions. Accuracy after removing or perturbing $k$ concept channels (RN/50).
Dataset $k$ Top-$k$ Removal Random Removal Random Insertion
Breakfast 0 1.000 1.000 1.000
1 0.496 0.989 0.979
2 0.229 0.972 0.947
3 0.085 0.951 0.951
4 0.028 0.923 0.908
HMDB51 0 1.000 1.000 1.000
1 0.603 0.974 0.978
2 0.374 0.963 0.961
3 0.238 0.949 0.945
4 0.142 0.935 0.921
To illustrate the effect of concept manipulation, we revisit Figure 2. Ablating the most influential concept bow by zeroing its activations changes the prediction from the correct class to run (logit 8.20→6.79). Removing windows 1–4, where the bow is handled, shifts the output to talk (logit 6.75). In addition, we evaluate three interventions: (i) top-$k$ concept removal, (ii) random concept removal, and (iii) random insertion of noise into concept activations ($\epsilon\sim\mathcal{N}(0,\,\sigma^{2})$ with $\sigma=0.5$). On the dataset level, Table 3 shows that accuracy declines sharply only when the most influential concepts are removed, while random removal or noise insertion cause minimal degradation. This indicates that MoTIF’s predictions rely on a small, semantically meaningful subset of concepts and remain robust under noise [^2] Note that $k=0$ corresponds to perfect accuracy, since all values are normalized relative to MoTIF’s baseline predictions..
4 Related Work
MoTIF lies at the intersection of three research areas: concept bottleneck models, temporal modeling, and video classification with vision–language backbones. At its core, MoTIF belongs to the family of concept bottleneck models (CBMs), which enforce intermediate, interpretable representations, while incorporating transformer-based sequence modeling that situates it within video classification and activity recognition. Below, we review related work and outline how MoTIF extends these directions.
Concept bottleneck models. CBMs (Koh et al., 2020) predict concepts as intermediate features before the final class prediction (Havasi et al., 2022; Chauhan et al., 2023; Prasse et al., 2025; Yang et al., 2023; Sawada and Nakamura, 2022). Recent research has focused on automatic concept discovery: DCLIP aligns data with CLIP’s vision–language space (Menon and Vondrick, 2023), LaBo queries large language models for diverse candidate concepts (Yang et al., 2023), and DCBM leverages segmentation foundation models to extract object- and part-level concepts (Prasse et al., 2025). Other works have extended this idea to different modalities. For instance, Ismail et al. (2025) adapt the bottleneck principle to protein design, where interpretable biochemical features form the concept space. Wu et al. (2022) extend CBMs to medical time series data, demonstrating that clinically meaningful temporal features can act as concepts. More recently, Sun et al. (2025) apply the CBM framework to language models, where intermediate concepts correspond to interpretable linguistic or semantic units.
While CBMs have been widely explored in the image domain, their extension to sequential data has seen little attention. Jeyakumar et al. (2022) propose extracting concepts from video descriptions and using them as inputs to a CBM for video classification, but their approach operates on a global level and thus produces concept activations that are not tied to specific parts of a video. Lee et al. (2025a) propose DANCE, which disentangles video concepts by combining pose sequences with textual concepts, providing users with a two-sided comparison between textual concepts and motion-dynamic concepts. In contrast, MoTIF generalizes the CBM principle from static images to arbitrary-length temporal sequences and focuses on agentic discovery of purely semantic textual concepts grounded in video data.
Video classification and action recognition. Video classification not only assigns a label to an entire sequence but, in the case of action recognition, also requires identifying the actions occurring within it (Pareek and Thakkar, 2021). Both tasks have progressed from CNN-based architectures (Tran et al., 2015; Lin et al., 2019; Liu et al., 2021) to transformer-based (Wang et al., 2024a) models such as TimeSformer (Bertasius et al., 2021) and VideoMAE (Tong et al., 2022), which capture long-range dependencies and leverage large-scale self-supervised pre-training. MoTIF integrates these approaches by adapting a transformer block for temporal modeling while incorporating a 1D-CNN to enable per-channel computations.
Temporal concept modeling. While concept-based explanations have been widely studied in the image domain (Knab et al., 2025a; Ghorbani et al., 2019), only a few works explore temporal dynamics (Gulshad et al., 2023; Kowal et al., 2024). Ji et al. (2023) propose a spatio-temporal concept framework to analyze representations in 3D ConvNets. PCBEAR (Lee et al., 2025b) introduces static and dynamic pose concepts for action recognition, and Saha et al. (2024) extend TCAV to videos by computing concept importance scores across sequences. 3D-ACE explains video models post hoc using user-defined concepts or supervoxels (Ghorbani et al., 2019). In contrast, MoTIF integrates concept-specific temporal attention directly into the predictive model, enabling reasoning over evolving concepts rather than treating them as fixed inputs or external explanations.
5 Discussion
MoTIF introduces a transformer-based concept bottleneck architecture for temporal data. The core architectural contribution of MoTIF is a temporal concept reasoning module that operates on sequences of interpretable concept activations. By combining per-concept temporal self-attention with transformer-style processing, MoTIF models when individual concepts recur and how their temporal patterns contribute to video-level predictions, while preserving explicit concept attribution. Across datasets, this design consistently outperforms zero-shot classification, global CBMs that average over windows, and prior video CBMs, demonstrating that transformer-based temporal reasoning can be integrated into concept bottleneck models without sacrificing interpretability.
Temporal concept quality is decisive for performance. While the MoTIF architecture provides the capacity for temporal reasoning, its effectiveness depends critically on the availability of concepts that reflect actions and interactions over time. Agentic concept discovery substantially improves performance—most notably on SSv2—by supplying object- and action-centric concepts that are temporally expressive. Together, these findings indicate that concept discovery and architectural advances are complementary: MoTIF provides the temporal reasoning backbone, while agent-generated concepts supply the temporally expressive evidence needed to handle challenging datasets.
MoTIF is modular and compatible with modern video backbones. We demonstrate that MoTIF can be paired with a range of embedding backbones, including CLIP and video-adapted Perception Encoders, with consistent gains across architectures. MoTIF-ST extends the framework to space–time transformers, showing that the proposed concept reasoning module is not restricted to a single transformer design. This modularity enables future integration with stronger video representations (Wang et al., 2024b) or alternative concept sources (Prasse et al., 2025; Lee et al., 2025b), while retaining interpretable temporal reasoning.
Interpretability is central to both explanation and model analysis. MoTIF’s three-level explanation interface—global, local, and temporal—exposes how individual concepts contribute across time without the entanglement introduced by standard attention mechanisms. This enables errors to be traced back to missing or ambiguous temporal concepts rather than opaque feature interactions, guiding refinement of the concept set and model behavior while maintaining competitive performance.
Challenges and future work. Selecting an appropriate temporal window size remains an important design consideration, as actions span variable durations and temporal granularity. A key next step for the community is to evaluate adaptive and multi-scale windowing strategies to improve robustness without sacrificing interpretability. Improving concept discovery—particularly for abstract interactions—is a high-impact next step: the substantial gains from agentic concepts on SSv2 indicate that richer concept sets yield larger benefits than architectural scaling alone. Finally, allowing cross-concept interactions increases expressivity but breaks faithful attribution; MoTIF makes this trade-off explicit rather than implicit by enforcing concept-level temporal isolation, while future work may explore controlled relaxations of this constraint.
6 Conclusion
We introduced MoTIF, a transformer-based concept bottleneck framework for video that combines temporal concept reasoning with agentic concept discovery. MoTIF employs diagonal temporal attention to track individual concepts over time, enabling faithful attribution of when and which concepts drive predictions while exposing a controllable accuracy–interpretability trade-off, where cross-concept attention can recover performance on demanding datasets at the cost of transparency. Our agentic discovery module enables the unsupervised discovery of temporally expressive object- and action-centric concepts directly from video, substantially narrowing the performance gap between interpretable models and black-box video transformers on challenging benchmarks. Overall, MoTIF provides a principled and scalable framework for learning, analyzing, and explaining temporal concept representations, enabling interpretable video understanding with competitive performance.
Acknowledgements
This research was supported in part by the German Federal Ministry for Economic Affairs and Climate Action of Germany (BMWK) and in part by the German Federal Ministry for Research, Technology, and Space (BMFTR), and was partially conducted during an internship at Ramblr.ai.
Appendix A Implementation Details
A.1 Hyperparameter Settings
Unless noted otherwise, all experiments in the main paper follow these default hyperparameters:
  • Training. 100 epochs, batch size 32, AdamW with learning rate $10^{-3}$ and weight decay $10^{-2}$.
  • Pooling. Log-sum-exp pooling with fixed temperature $\tau=1.0$. Although LSE permits tuning of sharpness, we kept $\tau$ constant for unbiased comparisons.
  • Regularization. Both the $\ell_{1}$ penalty on classifier weights and the activation sparsity penalty are set to $10^{-3}$.
  • Architecture. One Transformer layer with per-channel (diagonal) attention; classifier weights constrained to be nonnegative. As stated in Section 2, the diagonal attention faces the trade-off: interpretability through concept isolation versus representational power. Thus, additional depth yields diminishing returns under strict concept isolation, motivating systematic study of selectively enabling cross-concept interactions.
  • Data. Window size of 16 frames per temporal unit. For HMDB51 and Breakfast we report results on split s1, for UCF101 on testlist01, and for HMDB51 on split1. Class weighting is applied to mitigate imbalance.
Dataset-specific deviations: for SSv2 and UCF101 we use learning rate $0^{-4}, 1\ell_{1}$ penalty $10^{-4}$, and sparsity penalty $10^{-4}$; for SSv2, the non-negativity constraint in AdamW is disabled and batch size 128. HMDB51 uses a shorter window size of 8. For Breakfast, the window size is increased to 32 to account for longer videos.
All experiments were run with a random seed of 42 for reproducibility. Section C further analyzes the sensitivity to random seed choice.
These values serve as the baseline configuration; modifications for ablation studies are detailed in Section 3.2.2.
A.2 Computation Time and Complexity
Table 4 reports GPU memory (MB), epoch time (s) and throughput (samples/s) for HMDB51 and Breakfast using the two ablated clip backbones. Results are shown for diagonal attention enabled (On) and disabled (Off).
Table 4: Complexity overview. GPU memory, epoch time and throughput for RN/50 and B32 backbones with diagonal attention On/Off.
Dataset Backbone Setting GPU memory (MB) Epoch time (s) Samples / s
HMDB51 RN/50 On 1723 0.91 4084
Off 517 0.97 4063
B32 On 2289 0.90 4108
Off 1064 0.95 4071
Breakfast RN/50 On 10634 4.26 454
Off 736 0.51 3461
B32 On 11463 4.28 445
Off 1565 0.53 3445
Average $T$ (train set) HMDB51: 12.5    Breakfast: 65.4
For short videos (small $T$) epoch times remain nearly unchanged while GPU memory differs substantially. For longer sequences, as in Breakfast, disabling diagonal attention reduces memory use considerably; the runtime and memory gap grows with sequence length since complexity increases with $T$ (see Sec. 2) if $H\ll C$. The runtime across backbones is similar in magnitude; for example the best-performing perception encoder in MoTIF required 4.55 s. Video embedding time is excluded because embeddings are reusable across models and vary with the embedding backbone.
A.3 Backbone Integration
Image-based. Image–text models such as CLIP (Radford et al., 2021) and SigLIP (Zhai et al., 2023) embed single images rather than videos. To adapt them, we divide each video into windows of $F$ frames and randomly select one frame per window. This frame is embedded and serves as the window representation. Random sampling ensures variation in which part of the window is captured.
Video-based. The Perception Encoder (Bolya et al., 2025) is explicitly tuned to aggregate information across frames, producing a pooled embedding for each window. Unlike the image-based backbones, we therefore use multiple frames per window. In our experiments, we consistently used eight frames per window, except for HMDB51 where the window size itself was eight, so only four frames were sampled.
A.4 Space-Time Attention Extension
While the standard MoTIF architecture applies temporal attention independently for each concept, we extend it with a factorized space-time attention mechanism (Bertasius et al., 2021) that enables concept interactions while preserving interpretability.
The CBMTransformerST architecture factorizes attention into two sequential components: spatial attention across concepts at each time step, followed by the original per-channel temporal attention.
PerTimeSpatialBlock. This module computes attention across concepts at each time step independently. Given input features $X\in\mathbb{R}^{B\times T\times C}$, spatial attention produces attention scores $W_{s}\in\mathbb{R}^{B\times T\times C\times C}$ where $W_{s}[b,t,i,j]$ represents the attention weight from concept $i$ to concept $j$ at time step $t$.
Unlike temporal attention, which maintains concept isolation, spatial attention allows concepts to interact within each temporal frame. To preserve concept interpretability while enabling controlled spatial interaction, we employ three identity-preserving mechanisms:
  • Identity bias: We add a bias term $\alpha_{I}=1.0$ to the diagonal of the attention score matrix, encouraging self-attention and reducing cross-concept mixing.
  • Spatial gating: The spatial attention output is scaled by a gating factor $\beta_{s}\in[0,1]$ before being added to the residual connection: $X^{\prime}=X+\beta_{s}\cdot\text{SpatialAttn}(X)$. With $\beta_{s}=0.1$ (default), the spatial mixing contributes 10% to the output, preserving 90% of the original concept identity.
  • Per-channel FFN: The feed-forward network in the spatial block uses per-channel convolutions (grouped by $C$), ensuring no cross-concept mixing occurs in the FFN, matching the design of the temporal block.
These mechanisms ensure that concepts remain separable and interpretable: the spatial mixing is controlled and identity-preserving, the subsequent per-channel temporal attention maintains diagonal structure, and the final concept activations $Z_{t,c}$ and spatial attention maps $W_{s}$ enable attribution to individual concepts while revealing their interactions.
SpaceTimeBlock. The block applies spatial attention first to enable concept interactions at each time step, followed by per-channel temporal attention that maintains diagonal structure.
This factorization reduces computational complexity from $\mathcal{O}(T^{2}C^{2})$ for full space-time attention to $\mathcal{O}(TC^{2}+CT^{2})$, while producing separate interpretable attention maps for spatial ($W_{s}\in\mathbb{R}^{B\times T\times C\times C}$) and temporal ($W_{t}\in\mathbb{R}^{B\times C\times T\times T}$) components.
Exemplary Explanations. MoTIF with CBMTransformerST (MoTIF-ST) applies two sequential attention mechanisms—spatial followed by temporal—yielding two complementary attention matrices for interpretation.
  • The spatial attention block captures concept–concept interactions within each frame. It indicates which concepts influence a given concept and highlights additional concepts that are relevant at that moment. This provides a structured view of how concepts relate to each other spatially.
  • The temporal attention block captures when a concept becomes important. It identifies the time steps at which a specific concept contributes most to the final prediction, thereby revealing the temporal structure of the activity.
Figure 6 shows two examples from Breakfast and UCF101, including their attention matrices and the corresponding explanations.
Figure 6: MoTIF-ST explanations. Correctly classified examples from Breakfast and UCF101, illustrating MoTIF with a space–time transformer (ViT-L/14) and the corresponding temporal and spatial attention matrices.
Appendix B Additional Experiments
B.1 Where motifs help to understand Misclassifications
In Figure 7, we illustrate representative failure cases for each dataset using models trained with ViT-L/14. All corresponding videos, including the full set of temporal concepts across all windows, are provided in the supplementary material.
Figure 7: MoTIF explanations. Example videos from all datasets with incorrect classifications.
The first example, from Pirates of the Caribbean, shows Barbossa eating an apple before throwing it away. Our model predicts eat, while the ground truth is throw. Global and local concept attributions reveal that the concept eat, triggered by the apple, was most strongly activated. The corresponding attention map illustrates that early query frames attend to early key frames, indicating that the model anchors the decision on the moment where the apple is clearly visible and eaten. This concentrated attention explains why the model emphasizes eat over throw.
In the second example from UCF101, MoTIF detects correct concepts, such as basketball and scoreboard. However, these were insufficient to discriminate between the actions basketball dunk and basketball, leading to misclassification. Since the background frame remains nearly constant and only the players on the court are moving, the attention map for scoreboard shows a uniform distribution across time steps, with no clear temporal anchors. This indicates that the concept is consistently present.
The third example, from SSv2, highlights the dataset’s inherent difficulty. Unlike the correctly classified case in Figure 13, MoTIF predicts tearing sth just a little bit instead of the ground truth tearing sth into two pieces. Concept activations focus primarily on hand movements, such as crumple, together with the paper. The attention map shows that the concept crumple receives strongest attention from the later key frames (u=2, u=3) across several query times while other frames receive lower, more diffuse weights. This diffuse attribution explains why MoTIF captures the general action but fails to resolve the fine-grained distinction required by SSv2.
Finally, in the Breakfast dataset, MoTIF correctly identifies and attends to concepts relevant for egg. Yet, the dataset contains two distinct egg-related actions, friedegg and scrambledegg, which leads to misclassification. Nevertheless, all identified concepts correspond to actions of a person cooking with eggs. Furthermore, the attention map for grill reveals a broad and diffuse distribution across many time steps, indicating that the concept is persistently active throughout the sequence rather than concentrated in a few decisive moments. This persistent but unspecific attention explains why MoTIF captures the general presence of egg-related cooking but fails to distinguish between the two fine-grained classes.
Although MoTIF does not always predict the correct class, its structure makes the reasoning process transparent by decomposing decisions into concepts and their temporal interactions. In this work, our emphasis was on the architecture rather than on designing the most suitable concept sets. Follow-up experiments should evaluate stronger video-specific concept extraction pipelines to improve performance while retaining MoTIF’s interpretability.
B.2 Diagonal vs. Full Attention
We illustrate why full attention produces non-interpretable results by revisiting previously shown examples. In Figure 2, replacing diagonal attention with full attention shifts the most important concepts to frown (activation 4.796) and face (1.282), while all others fall below 0.01 and are omitted. Although these concepts appear as locally relevant, they do not correspond to the depicted action (class bow). This indicates that concept mixing occurs: full attention entangles channels, creates arbitrary concepts that are not visually apparent and thus undermine interpretability.
B.3 Interpreting Temporal Dependencies
An attention map visualizes how a concept channel distributes its focus across time. Each entry $W_{c,t,u}$ encodes the attention weight between query time step $t$ (vertical axis) and key/value time step $u$ (horizontal axis). A bright cell at position $(t,u)$ indicates that the activation of concept $c$ at time $t$ strongly attends to the representation of the same concept at time $u$. Diagonal patterns suggest that the concept mainly attends to itself at the same or nearby frames, while vertical stripes show that many query frames refer back to the same key frame, indicating the presence of a temporal anchor. In contrast, diffuse or uniform maps imply that the concept is expressed consistently across time rather than being tied to specific moments. Thus, by inspecting these maps, one can infer whether a concept is localized, persistent, or temporally linked to particular frames within the sequence.
B.4 Synthetic Temporal Data
The synthetic experiment is designed to isolate temporal reasoning by removing all appearance cues. Each “video” is a sequence of length-N (matching the Breakfast distribution), where each frame is a 1024-dimensional vector. We define five temporal classes: (1) ascending, (2) descending, (3) U-shape, (4) inverted-U, and (5) periodic oscillation. Class identity is determined only by the temporal trajectory of 30% of the dimensions; the remaining dimensions are filled with noise. Thus, the task can only be solved by learning the temporal order—static appearance information is uninformative by design.
To examine whether the model can learn interpretable temporal structure, we provide 37 artificial temporal pattern concepts. Each concept represents a prototypical analytic temporal signal (ascending, descending, S-curve, exponential growth/decay, chirp, sawtooth, step-up/down, periodic-fast, oscillating-decay, double-peak, plateau, etc.). All embeddings encode temporal shape only, without semantics.
The training setup matches our real pipelines, allowing a direct test of MoTIF’s temporal inductive biases. MoTIF achieves 86.97% accuracy on the original (ordered) sequences but only 21.06% when frames are randomly permuted (chance 20%). This 4.1× drop shows that MoTIF relies on temporal ordering rather than static frame statistics. An image-level CBM baseline (“Global CBM”), which collapses all frames into a single vector, reaches only 35.5%, confirming that methods without temporal modeling cannot solve this task.
Appendix C Additional Ablations
(a) Non-negativity. Test accuracy with and without enforcing non-negativity.
(b) Weight decay. Test accuracy across different weight decays.
Figure 8: Architectural choices. Effects of non-negativity and weight decay.
For completeness, we report further ablation studies that complement the main paper (Section 3.2.2).
Nonnegativity. The non-negativity constraint on classifier weights improves interpretability by ensuring class predictions are explained solely by positive concept evidence (Zhang et al., 2021). We enforce non-negativity of classifier weights $W$ by projection after each update. Results are shown in Figure 8(a). This constraint tends to cause minor accuracy reductions for RN/50 on both datasets, while for B/32 on Breakfast we observe a slight improvement. Given the small fluctuations inherent in training, these effects should be interpreted as indicative rather than strictly conclusive. Overall, the effect on test accuracy is negligible. We therefore enable it for all experiments except SSv2, where modeling negative concepts is required to capture fine-grained actions.
Weight decay. Figure 8(b) reports test accuracy across different weight decay values for the AdamW optimizer. Performance remains largely stable, similar to the non-negativity constraint, except for Breakfast using B/32. We therefore fix the weight decay to $10^{-2}$ for all experiments.
Figure 9: Window size influence. Test accuracy across different window sizes.
Per-concept affine transformation. We optionally insert a per-concept affine transformation between the per-channel temporal self-attention and the classification head. While not strictly required, this block sharpens activations by rescaling and shifting each concept dimension before the nonlinearity. Figure 10 shows the effect on Breakfast and HMDB51: enabling the affine transformation yields a slight increase in test accuracy and a consistent reduction in both logits and concept entropy, indicating sharper and more decisive concept activations, especially on the Breakfast dataset.
Figure 10: Effect of the per-concept affine transformation. Accuracy improves marginally, while entropy in both logits and concept activations decreases, suggesting that the affine block stabilizes and sharpens the CBM’s internal representations.
Classifier sparsity. The sparsity penalty on classifier weights has a pronounced impact on test accuracy, as shown in Figure 11(a). Larger values of $\lambda_{\ell_{1}}$ consistently reduce accuracy. We therefore set $\lambda_{\ell_{1}}=10^{-3}$ in most experiments, as it offers a reasonable trade-off between regularization and performance.
Activation sparsity. The activation sparsity penalty shows little variation in test accuracy (see Figure 11(b)) across different values of $\lambda_{\mathrm{sparse}}$, except for Breakfast. We attribute this to the comparatively long video sequences in that dataset. Accordingly, we use $\lambda_{\mathrm{sparse}}=10^{-3}$ for all experiments, except for Breakfast, where we set it to $10^{-4}$.
Learning rate. The learning rate has a strong effect on test accuracy, as shown in Figure 11(c). For Breakfast and HMDB51, we set it to $10^{-3}$, while for UCF101 and SSv2, we use $10^{-4}$, which yielded better performance, an effect was not seen in these ablations.
Window size. We evaluate MoTIF with varying temporal input lengths to study robustness to sequence duration and efficiency trade-offs. While increasing the number of frames provides more temporal context, it also raises memory requirements and may not yield consistent accuracy gains. For comparability, we fix the batch size to $8$ across all ablations, ensuring that changes in performance are solely attributable to window size rather than training dynamics. The results in Fig. 9 highlight that optimal window size depends on both dataset characteristics and backbone choice.
(a) Classifier sparsity. Test accuracy with different $\lambda_{\ell_{1}}$.
(b) Activation sparsity. Test accuracy across different $\lambda_{\mathrm{sparse}}$.
(c) Learning rate. Test accuracy with different learning rates.
Figure 11: Architectural choices. Effects of classifier sparsity, activation sparsity, and learning rate.
Figure 12: Random seed. Box plot of ablated datasets with five different seeds.
Random seed. Since most experiments were run with a fixed seed of 42 for reproducibility, we additionally ablate the effect of varying the random seed (39,40,41,42,43) for both MoTIF and the global CBM. As Figure 12 illustrates, the influence of the seed depends on the dataset. For HMDB51 the effect is negligible, whereas for Breakfast — particularly with the RN/50 backbone — we observe noticeably higher variance. Nevertheless, MoTIF consistently outperforms the global CBM. Moreover, the figure indicates that the reported numbers in Table 1 are conservative: even higher scores are achievable, but we deliberately refrain from reporting maxima and instead provide a representative overview.
Additionally, since embeddings depend on the random choice of a representative image per window (previously fixed at seed 42), we re-embedded each dataset with five seeds. For each embedding variant we trained and evaluated MoTIF, the global CBM , and the zero-shot runs using a fixed training seed of 42 to isolate the effect of image-selection randomness. Table 5 reports the mean and standard deviation across embeddings (seeds 39–43). Std values are small overall; the largest observed variability is for Breakfast with RN/50 (std = 1.1).
Table 5: Random seed on embeddings. Mean and standard deviation (%) for methods on HMDB51 and Breakfast with two backbones.
Dataset (backbone) Method Mean (%) Std (%)
HMDB51 (RN/50) MoTIF 63.8 0.3
Global CBM 60.2 0.3
Zero Shot 29.4 0.2
HMDB51 (B32) MoTIF 66.5 0.3
Global CBM 61.0 0.6
Zero Shot 38.7 0.3
Breakfast (RN/50) MoTIF 53.2 1.1
Global CBM 41.1 1.1
Zero Shot 18.1 0.7
Breakfast (B32) MoTIF 58.3 0.8
Global CBM 43.9 0.6
Zero Shot 25.8 0.2
C.1 Accuracies Testsplits
Tables 68 report Top-1 accuracies (%) on the test splits. Mean and standard deviation are computed across the shown splits (Breakfast: 4 splits; HMDB51 and UCF: 3 splits). For each backbone we list the MoTIF result (bold), the corresponding global CBM, and the zero-shot baseline. The reported standard deviation quantifies variability between test splits.
Table 6: Top-1 accuracies (%) on Breakfast test splits. Mean and standard deviation across splits.
Model s1 s2 s3 s4 Mean Std
CLIP RN/50 (MoTIF) 55.6 47.4 47.0 61.2 52.8 6.9
Global CBM (RN/50) 43.3 25.7 32.5 44.4 36.5 9.0
Zero-shot (RN/50) 17.3 19.2 16.0 21.9 18.6 2.6
CLIP B/32 (MoTIF) 56.3 44.9 51.3 61.0 53.4 6.9
Global CBM (B/32) 42.6 27.5 31.6 47.0 37.2 9.1
Zero-shot (B/32) 24.6 19.8 22.0 26.4 23.2 2.9
CLIP L/14 (MoTIF) 71.5 62.3 66.7 76.8 69.3 6.2
CLIP L/14 (MoTIF-ST) 73.2 64.8 65.0 80.8 73.5 12.2
Global CBM (L/14) 61.3 44.7 48.7 66.4 55.3 10.2
Zero-shot (L/14) 32.4 28.3 26.5 37.0 31.1 4.7
SigLIP L/14 (MoTIF) 76.1 62.1 73.1 82.7 73.5 8.6
Global CBM (SigLIP L/14) 59.9 44.7 53.3 70.6 57.1 10.9
Zero-shot (SigLIP L/14) 28.9 18.0 20.9 26.6 23.6 5.0
PE L/14 (MoTIF) 87.3 74.7 83.1 89.4 83.6 6.5
PE L/14 (MoTIF-ST) 86.2 74.7 82.3 91.2 83.6 7.0
Global CBM (PE L/14) 81.0 58.7 72.0 79.9 72.9 10.3
Zero-shot (PE L/14) 40.5 36.6 36.8 51.6 41.4 7.0
Table 7: Top-1 accuracies (%) on HMDB51 test splits. Mean and standard deviation across splits.
Model s1 s2 s3 Mean Std
CLIP RN/50 (MoTIF) 64.1 62.3 62.1 62.8 1.1
Global CBM (RN/50) 58.6 60.1 59.2 59.3 0.8
Zero-shot (RN/50) 29.3 30.1 30.1 29.8 0.5
CLIP B/32 (MoTIF) 65.9 66.8 63.3 65.3 1.8
Global CBM (B/32) 62.0 63.0 59.8 61.6 1.6
Zero-shot (B/32) 38.4 37.9 38.0 38.1 0.3
CLIP L/14 (MoTIF) 73.8 73.9 72.2 73.3 1.0
CLIP L/14 (MoTIF-ST) 75.5 75.3 73.6 74.8 1.0
Global CBM (L/14) 68.5 68.8 67.9 68.4 0.5
Zero-shot (L/14) 45.8 45.6 45.6 45.7 0.1
SigLIP L/14 (MoTIF) 74.8 74.4 70.4 73.2 2.4
Global CBM (SigLIP L/14) 66.3 66.0 62.6 65.0 2.1
Zero-shot (SigLIP L/14) 48.4 50.0 49.5 49.3 0.8
PE L/14 (MoTIF) 79.9 79.3 79.6 79.6 0.3
PE L/14 (MoTIF-ST) 79.9 80.0 78.8 79.6 0.7
Global CBM (PE L/14) 74.0 75.0 74.1 74.4 0.6
Zero-shot (PE L/14) 56.7 57.3 56.2 56.7 0.6
Table 8: Top-1 accuracies (%) on UCF test splits. Mean and standard deviation across splits.
Model s1 s2 s3 Mean Std
CLIP RN/50 (MoTIF) 82.4 82.5 83.4 82.8 0.6
Global CBM (RN/50) 80.7 80.0 79.3 80.0 0.7
Zero-shot (RN/50) 56.5 56.9 58.3 57.2 0.9
CLIP B/32 (MoTIF) 84.4 85.7 86.7 85.6 1.2
Global CBM (B/32) 82.2 82.5 83.6 82.8 0.7
Zero-shot (B/32) 59.4 60.2 60.1 59.9 0.4
CLIP L/14 (MoTIF) 92.4 93.7 93.6 93.2 0.7
CLIP L/14 (MoTIF-ST) 92.9 94.4 94.0 93.8 0.8
Global CBM (L/14) 88.8 91.0 90.3 90.0 1.1
Zero-shot (L/14) 71.1 70.6 70.1 70.6 0.5
SigLIP L/14 (MoTIF) 93.3 93.8 94.9 94.0 0.8
Global CBM (SigLIP L/14) 90.0 91.0 90.5 90.5 0.5
Zero-shot (SigLIP L/14) 80.0 81.9 79.2 80.4 1.4
PE L/14 (MoTIF) 94.6 95.7 95.8 95.4 0.7
PE L/14 (MoTIF-ST) 95.7 96.4 96.9 96.3 0.6
Global CBM (PE L/14) 94.6 93.9 95.0 94.5 0.6
Zero-shot (PE L/14) 73.8 75.6 74.4 74.6 0.9
C.2 Performance on the Ablation Concept Set
We additionally evaluate all methods using the non-agentic concept sets generated via an LLM, as described in Section 2.1. The corresponding results are reported in Table 9. Overall, both Global CBM and MoTIF exhibit slightly reduced performance compared to the agentic setting, which we attribute to the smaller and less complete concept set, limiting the amount of task-relevant information captured in the bottleneck.
Table 9: Performance comparison (% Top-1 accuracy). Mean $\pm$ standard deviation on train-test splits on Breakfast Actions, HMDB51, UCF101, and SSv2 with different CLIP-based backbones with the non-agentic concept set. We report seconds per training epoch next to the accuracy scores for all MoTIF variants.
Method Breakfast HMDB51 UCF101 SSv2
Zero-shot
CLIP-RN/50 (Radford et al., 2021) 18.6 $\pm$ 2.6 29.8 $\pm$ 0.5 57.2 $\pm$ 0.9 0.8
CLIP-ViT-B/32 (Radford et al., 2021) 23.2 $\pm$ 2.9 38.1 $\pm$ 0.3 59.9 $\pm$ 0.4 0.9
CLIP-ViT-L/14 (Radford et al., 2021) 31.1 $\pm$ 4.7 45.7 $\pm$ 0.1 70.6 $\pm$ 0.5 0.9
SigLIP-L/14 (Zhai et al., 2023) 23.6 $\pm$ 5.0 49.3 $\pm$ 0.8 80.4 $\pm$ 1.4 1.3
PE-L/14 (Bolya et al., 2025) 41.4 $\pm$ 7.0 56.7 $\pm$ 0.6 74.6 $\pm$ 0.9 2.2
Global CBM
CLIP-RN/50 (Radford et al., 2021) 36.5 $\pm$ 9.0 59.3 $\pm$ 0.8 80.0 $\pm$ 0.7 13.7
CLIP-ViT-B/32 (Radford et al., 2021) 37.2 $\pm$ 9.1 61.6 $\pm$ 1.6 82.8 $\pm$ 0.7 15.2
CLIP-ViT-L/14 (Radford et al., 2021) 55.3 $\pm$ 10.2 68.4 $\pm$ 0.5 90.0 $\pm$ 1.1 18.1
SigLIP-L/14 (Zhai et al., 2023) 57.1 $\pm$ 10.9 65.0 $\pm$ 2.1 90.5 $\pm$ 0.5 19.6
PE-L/14 (Bolya et al., 2025) 72.9 $\pm$ 10.3 74.4 $\pm$ 0.6 94.5 $\pm$ 0.6 25.5
MoTIF (ours)
MoTIF (RN/50) 52.8 $\pm$ 6.9 (4.2) 62.8 $\pm$ 1.1 (0.9) 82.8 $\pm$ 0.6 (1.5) 16.0 (10.0)
MoTIF (ViT-B/32) 53.4 $\pm$ 6.9 (4.2) 65.3 $\pm$ 1.8 (0.9) 85.6 $\pm$ 1.2 (1.5) 17.5 (9.9)
MoTIF (ViT-L/14) 69.3 $\pm$ 6.2 (4.3) 73.3 $\pm$ 1.0 (0.8) 93.2 $\pm$ 0.7 (1.5) 20.4 (10.1)
MoTIF-ST (ViT-L/14) 71.1 $\pm$ 7.7 (7.2) 74.8 $\pm$ 1.0 (1.8) 93.8 $\pm$ 0.9 (3.3) 23.9 (26.8)
MoTIF (SigLIP-L/14) 73.5 $\pm$ 8.6 (4.2) 73.2 $\pm$ 2.4 (0.8) 94.0 $\pm$ 0.8 (1.5) 22.4 (9.8)
[…]
MoTIF-ST (PE-L/14) 84.1 $\pm$ 6.4 (7.3) 79.6 $\pm$ 0.7 (1.8) 96.3 $\pm$ 0.6 (3.3) 35.1 (26.7)
Non-interpretable video models
TSM (Lin et al., 2019) 59.111footnotemark: 1 73.5 95.9 61.7
No frame left behind (Liu et al., 2021) 62.011footnotemark: 1 73.411footnotemark: 1 96.411footnotemark: 1 62.711footnotemark: 1
VideoMAE V2 (Wang et al., 2023) 88.1 99.6 76.8
Appendix D Algorithms
This section summarizes MoTIF’s procedures for training, test-time inference, and explanation. We adopt the notation from the main text: per-window concept activations $Z_{t,c}$, per-time logits $\ell_{t}$, and log-sum-exp (LSE) pooling with temperature $\tau$ and optional mask $m_{t}$.
Algorithm 1 Training MoTIF (Moving Temporal Interpretable Framework)
1:Input: $\{(X^{(n)},y^{(n)})\}_{n=1}^{N}$, concept bank $\mathcal{C}$
2: Initialize Transformer parameters; affine $(\gamma,\delta)$; classifier $(W,b)$
3:for each epoch do
4:  for each batch $(x,y)$ do
5:   Temporal modeling: per-channel temporal self-attention $\to X^{(L)}_{t,c}$
6:   Affine & nonnegativity: $Z_{t,c}\leftarrow\mathrm{Softplus}(\gamma_{c}X^{(L)}_{t,c}+\delta_{c})$
7:   Classification: $\ell_{t}\leftarrow WZ_{t,:}+b$
8:   Pooling: $\hat{\ell}\leftarrow\mathrm{LSE}_{\tau}(\{\ell_{t}\},m)$
9:   Loss: $\mathcal{L}\leftarrow\mathrm{CE}(\hat{\ell},y)+\lambda_{\ell_{1}}\|W\|_{1}+\lambda_{\mathrm{sparse}}\frac{1}{(\sum_{t}m_{t})C}\sum_{t,c}m_{t}|Z_{t,c}|$
10:   Update all parameters with AdamW
11:   Optional: enforce nonnegativity $W\leftarrow\max(W,0)$
12:  end for
13:end for
14:Output: trained MoTIF
Description. The training loop builds concept activations from window embeddings, refines them with per-channel temporal attention, applies a nonnegative affine projection, and classifies with a linear head pooled over time via LSE. The objective combines cross-entropy with sparsity regularizers; $W$ can be projected to enforce nonnegativity.
Algorithm 2 Inference with MoTIF (test-time forward pass)
1:Input: video $x$, trained MoTIF
2: Compute concept activations from window embeddings
3: Apply per-channel temporal attention; affine + Softplus $\to$ nonnegative $Z_{t,c}$
4: Compute per-time logits $\ell_{t}\leftarrow W_{k}Z_{t,:}+b$
5: Aggregate with LSE pooling: $\hat{\ell}\leftarrow\mathrm{LSE}_{\tau}(\{\ell_{t}\},m)$
6: Predict $y^{*}\leftarrow\arg\max_{k}\hat{\ell}_{k}$
7:Output: predicted label $y^{*}$
Description. Inference is a single forward pass: per-window logits are pooled with LSE to produce video-level logits, whose argmax yields the prediction.
Algorithm 3 Explanation with MoTIF (global, local, temporal views)
1:Input: video $x$, class $k$, trained MoTIF
2: Run forward pass to obtain $Z_{t,c}$, $\ell_{t}$, and $\hat{\ell}$
3:Per-time contributions: $\mathbf{c}_{t}^{(k)}\leftarrow Z_{t,:}\odot W_{k,:}$
4:     $s_{t}^{(k)}\leftarrow\sum_{c}c_{t,c}^{(k)}+b_{k}$
5:Temporal importance: $\pi_{t}^{(k)}\leftarrow\mathrm{softmax}\!\big(s_{t}^{(k)}/\tau\big)$ over valid $t$ (mask $m_{t}$)
6:Global attribution: $\bar{\mathbf{c}}^{(k)}\leftarrow\sum_{t}\pi_{t}^{(k)}\,\mathbf{c}_{t}^{(k)}$
7:Output: $\big(\bar{\mathbf{c}}^{(k)},\{\pi_{t}^{(k)}\}_{t},\text{attention maps}\big)$
Description. Explanations decompose the prediction into per-concept contributions and reweight them by a time-importance distribution that mirrors LSE pooling. This yields three complementary views: (1) global concepts via $\bar{\mathbf{c}}^{(k)}$, (2) local concepts at decisive windows (large $\pi_{t}^{(k)}$), and (3) temporal dependencies from per-concept attention maps.
Algorithm 4 Agentic Concept Discovery for Video (objects + actions)
1:Input: training set $\mathcal{D}_{\text{train}}$, agent $\mathcal{A}$, windows $k$, frames/window $l$, max videos/class $n$, similarity threshold $\delta$
2:Output: filtered concept set $\mathcal{C}$
3: Initialize candidate pool $\mathcal{C}_{\text{raw}}\leftarrow[\,]$
4:for each class $y$ do
5:  Sample up to $n$ videos $\{x\}$ of class $y$ from $\mathcal{D}_{\text{train}}$
6:  for each video $x$ do
7:   Split $x$ into $k$ windows and sample $l$ frames per window
8:   Generate object and action concepts conditioned on $y$:
9:     $\mathcal{C}_{x}\leftarrow\mathcal{A}(\text{frames},y)$
10:   Append $\mathcal{C}_{x}$ to $\mathcal{C}_{\text{raw}}$
11:  end for
12:end for
13: Embed all concepts using text encoder $\psi(\cdot)$
14: Filter $\mathcal{C}_{\text{raw}}$ by cosine similarity threshold $\delta$ to obtain $\mathcal{C}$
15:return $\mathcal{C}$
Description. Agentic concept discovery decomposes videos into short temporal windows and leverages a vision–language agent to extract textual object and action concepts from sampled frames. Concepts are aggregated across windows and videos into a shared candidate pool and subsequently filtered using a similarity threshold to remove redundant or semantically overlapping entries. This results in a compact yet expressive concept set constructed in an unsupervised manner from the training data only, which serves as the bottleneck representation for MoTIF.
Visualized Pipeline. Figure 13 summarizes the MoTIF pipeline. Given an input video, temporal windows are first embedded using a frozen vision–language backbone. These embeddings are mapped to concept activations via cosine similarity to a predefined concept bank, yielding per-time, per-concept responses. MoTIF then applies per-channel temporal self-attention to model concept-specific dynamics independently over time. The resulting representations are passed through a nonnegative affine projection and a linear classifier, producing per-time logits that are aggregated into a video-level prediction using log-sum-exp pooling. Beyond prediction, the explicit concept bottleneck enables three complementary explanation views: global concept attributions aggregated over time, local concept activations at informative temporal segments, and temporal dependencies captured by the attention weights.
Figure 13: MoTIF pipeline. Videos are embedded with a vision–language backbone and mapped to concept activations via cosine similarity. Per-channel temporal self-attention models dynamics independently for each concept, followed by a nonnegative affine transformation and classification. MoTIF enables explanations across three views: global concepts, local concepts, and temporal dependencies. Sample frames from SSv2 (Materzynska et al., 2020) with MoTIF (ViT-L14).
Appendix E Concept Sets
E.1 Concept Discovery with Agentic Approach
Agent configuration. In our implementation, $A_{\mathrm{vlm}}$ receives a small set of representative frames from $w_{k}$ (uniformly sampled), along with a fixed prompt template that instructs the model to return (i) object/attribute concepts and (ii) action/interaction concepts as short noun or verb phrases. We use deterministic decoding (temperature 0) and request a bounded number of candidates per window (up to $M$ per category), which makes concept proposal reproducible and controls the size of the downstream concept bank. When class labels are available at training time, we optionally provide the video class as additional context to bias proposals toward class-discriminative concepts; at test time, no labels are provided.
Ablation on VLM Backbones. To study the impact of the underlying vision–language model, we repeat the concept discovery procedure using different Qwen 3 backbones (4B, 8B, and 30B parameters) Breakfast and HMDB51. Table 10 reports the resulting mean classification accuracy together with the average number of discovered concepts. Note that the number of concepts can vary across splits, as concepts are extracted solely from the training videos.
Table 10: Ablation on Qwen 3 backbones for agentic concept discovery. Mean Top-1 accuracy and average number of discovered concepts (train split only), evaluated with different CLIP vision backbones.
Qwen Vision Breakfast HMDB51
Backbone Acc. (%) #Concepts Acc. (%) #Concepts
Qwen 3 4B RN-50 55.2 338 66.2 793
ViT-B/32 53.3 240 68.6 618
Qwen 3 8B RN-50 54.5 330 66.8 872
ViT-B/32 54.1 238 68.6 662
Qwen 3 30B RN-50 55.1 258 66.2 657
ViT-B/32 52.7 193 68.5 511
Ablation on Similarity Threshold. We further ablate the effect of the similarity threshold used for concept filtering on Qwen 3 30B. While our main experiments use a threshold of 0.9, Table 11 reports results for thresholds of 0.8 and 0.7, together with the corresponding concept counts.
Table 11: Ablation on concept similarity threshold. Impact on concept count and downstream accuracy for different vision backbones.
Threshold Vision Backbone Breakfast HMDB51
Acc. (%) #Concepts Acc. (%) #Concepts
0.7 RN-50 39.5 24 56.1 51
ViT-B/32 29.9 11 49.4 27
0.8 RN-50 51.5 88 63.8 217
ViT-B/32 49.1 52 64.3 128
0.9 RN-50 55.1 258 66.2 657
ViT-B/32 52.7 193 68.5 511
For a threshold of 0.9, the number of filtered concepts ranges from 1,748–2,325 on SSv2, 481–1,217 on HMDB51, 186–366 on Breakfast, and 729–2,075 on UCF101. The large variance is primarily driven by the embedding model and the size of the training splits. Importantly, larger concept sets consistently lead to improved performance, indicating that increased concept diversity yields a more complete bottleneck representation.
E.2 LLM Concepts
In Table 12, we show the number and kind of concepts used for the construction of MoTIF for each dataset. The number and kind of concepts vary for each dataset, since we asked the LLM to create domain-specific concepts that are useful for the downstream classification task. As the tables indicate, the number of concepts is lower than the agentic based one.
Table 12: MoTIF concepts. The textual concepts utilized for all experiments which are listed in this paper.
Set Count Concepts
Breakfast 223 add, adjust, apple, arrange, assemble, avocado, bacon, bagel, bake, balance, banana, batter, beat, bin, blend, blender, blow, boil, bottle, bowl, bread, brew, brush, burner, butter, button, carry, carton, catch, cereal, chair, cheese, chop, cinnamon, clap, close, coffee, colander, comb, container, cook, cookbook, cool, core, counter, cover, crack, croissant, cucumber, cup, cupboard, cut, cuttingboard, detergent, dish, drag, drain, drizzle, drop, dry, egg, faucet, fill, flame, flip, fold, fork, freezer, fridge, froth, frown, fruit, fry, garbage, gesture, granola, grate, grater, grill, grind, ham, handle, heat, herb, hide, honey, hood, ice, ingredient, insert, jar, juice, kettle, knife, knob, knock, ladle, laugh, leftover, lid, mash, measure, measuringcup, measuringspoon, milk, mix, mug, napkin, nod, onion, open, orange, oven, ovenmitt, pack, package, pan, pantry, pastry, peel, peeler, pick, pinch, pit, place, plate, plug, point, poke, pour, preheat, press, pull, push, put, reach, recipe, recycle, release, remove, reveal, rinse, roll, rotate, sausage, scale, scoop, scramble, scrub, seal, serve, serving, set, shake, shave, sieve, sink, sip, sit, slice, slide, smile, snap, soap, socket, sort, spatula, spin, sponge, spoon, spread, sprinkle, squeeze, stack, stand, start, steam, steep, stir, stirrer, stool, stop, stove, strawberry, sugar, switch, syrup, table, take, tamp, tap, taste, tea, thermometer, throw, tie, tilt, timer, toast, tomato, tongs, toss, towel, tray, turn, twist, uncover, unfold, unscrew, unstack, untie, unwrap, warm, wash, waste, water, wave, whisk, wipe, wring, yogurt, zest, zip
UCF101 166 aim, archer, archery, arena, arrow, athlete, balance, ball, bar, barbell, baseball, basket, basketball, bat, beam, bicycle, block, bounce, bow, bowl, boxing, breakdance, canoe, cap, catch, clap, climb, club, coach, control, court, cricket, curl, dance, dancer, deadlift, dismount, dive, dodge, dribble, dumbbell, enter, field, fight, flip, floor, frisbee, gallop, gloves, goal, goalpost, grab, grapple, grind, gun, gym, gymnast, handstand, hang, helmet, hit, hockey, hook, hoop, horse, hurdle, ice, instrument, jab, jersey, jump, kayak, kick, ladder, lane, lift, mat, microphone, mount, music, net, netting, opponent, pad, paddle, parry, pass, pedal, perform, pitch, platform, player, pool, press, puck, pull, push, racket, rail, raise, referee, reins, release, reload, ride, ring, rope, row, rower, rugby, run, sand, scoreboard, serve, sheet, shoot, shooter, sit, skateboard, skateboarder, skater, ski, skip, skis, smash, snow, snowboard, snowboarder, soccer, spike, spin, splash, sprint, squat, stadium, stage, stand, start, steer, stick, stop, strike, surf, surfboard, surfer, swim, swimmer, swing, sword, target, teammate, throw, timer, track, trampoline, tuck, turn, uniform, uppercut, volleyball, walk, wall, water, wave, wrestle, yoga
HMDB51 150 apply, around, backward, balance, ball, baseball, basketball, bat, bend, bicycle, block, blow, bottle, bounce, bow, brake, brush, button, carry, cartwheel, catch, chair, chew, climb, close, comb, crawl, cross, crouch, cup, dismount, dive, door, down, drag, dribble, drink, drop, eat, enter, exit, face, fall, fight, finish, flip, float, frisbee, from, frown, gallop, grab, hair, hand, hands, handstand, hat, head, headstand, high, hit, hop, horse, hug, jacket, jog, juggle, jump, kick, kiss, knock, laugh, leap, left, leg, lie, lift, line, look, low, makeup, mount, mouth, nod, object, off, on, open, pedal, point, pull, punch, push, put, racket, reach, release, ride, right, roll, room, run, serve, shake, shave, shirt, shoelace, shoot, sing, sip, sit, skate, skateboard, ski, sled, sleep, slide, smile, snowboard, somersault, spin, sprint, stand, start, steer, stretch, surface, swim, swing, sword, take, talk, teeth, tennis, throw, tie, toss, touch, turn, untie, up, utensils, wake, walk, wash, wave, with, words, yawn, zip
SSv2 284 accelerate, apple, arm, assemble, background, backpack, bag, balance, ball, banana, bend, bite, blow, book, bottle, bottom, bounce, bow, bowl, box, break, broken, can, cap, carrot, carry, catch, chair, chew, chop, chopstick, clap, clean, click, climb, close, closeable, cold, connect, container, cough, cover, crawl, crouch, crumple, cry, cucumber, cup, cut, dance, decelerate, dirty, disassemble, disconnect, door, downward, drag, drag mouse, draw, drink, drinkable, drop, dry, durable, eat, edible, empty, erase, face, fall, fasten, fill, finger, fixed, flatten, flip, floor, fold, fork, fragile, frown, fruit, full, gather, get up, grape, hand, heavy, hide, hold, hop, hot, insert, inside, juggle, jump, key, keyboard, kneel, knife, knock, laptop, laugh, lean, left, lid, lift, light, lock, loosen, mix, mouse, nod, object, open, openable, orange, other, outside, paint, paper, peel, pen, pencil, person, phone, plate, plug, point, pour, pourable, press, pretend to balance, pretend to block, pretend to bow, pretend to catch, pretend to catch fish, pretend to clap, pretend to clean, pretend to climb, pretend to close, pretend to cook, pretend to dance, pretend to dodge, pretend to draw, pretend to dribble, pretend to drink, pretend to drive, pretend to eat, pretend to fall, pretend to fire gun, pretend to honk, pretend to hug, pretend to jump rope, pretend to kick, pretend to kiss, pretend to load gun, pretend to lock, pretend to look around, pretend to measure, pretend to open, pretend to paddle, pretend to paint, pretend to play drums, pretend to play guitar, pretend to play piano, pretend to point, pretend to pour, pretend to pull, pretend to punch, pretend to push, pretend to read, pretend to row, pretend to salute, pretend to scroll, pretend to search, pretend to serve, pretend to shake hands, pretend to shoot arrow, pretend to shoot basket, pretend to sing, pretend to sleep, pretend to steer, pretend to steer wheel, pretend to stir, pretend to swing bat, pretend to swipe, pretend to throw, pretend to throw ball, pretend to type, pretend to unlock, pretend to use controller, pretend to wake, pretend to wave, pretend to weigh, pretend to write, pull, push, remote, remove, reveal, right, roll, rollable, rotate, rough, run, scatter, scoop, scroll, separate, shake, shake head, shelf, shout, sip, sit, sleep, slice, slide, smell, smile, smooth, snap, sneeze, speak, spill, spillable, spin, spin dance, spit, spoon, sprinkle, sprint, squeezable, stack, stackable, stand, start, stir, stop, stretch, stumble, surface, swing, swipe, table, tap, taste, tear, throw, tie, tighten, tilt, tomato, top, topple, touch, toy, turn off, turn on, type, uncover, unfasten, unfold, unlock, unplug, unstack, untie, unwrap, upward, vegetable, wake, walk, wall, wave, wet, whisper, window, wrap, write, yawn, zoom in, zoom out
E.3 Concept Set Variance
Figure 14: Concept set influence. Distribution of test accuracy across five different concept sets. The two dots indicate outliers within the interquartile range.
Concept set influence. To investigate the influence of the concept proposal set on the performance of MoTIF, we prompt GPT-5 five times, each time requesting a distinct set of candidate concepts (with some natural overlap across runs). Our experiments demonstrate that given the same prompt to the LLM, concept set affects test accuracy only moderately, with consistent trends across datasets and backbones (Fig. 14). In line with prior work on concept bottleneck models, more dataset-specific concepts tend to increase accuracy (Rao et al., 2024; Prasse et al., 2025; Schrodi et al., 2025). However, our main motivation here is not to optimize the quality of the concept proposals, but rather to demonstrate that our framework works broadly and robustly across datasets, backbones, and varying concept sets. In addition, we note that kk—the number of concepts retained in the bottleneck—influences the achievable accuracy: very small $k$ restricts expressive power, while very large $k$ may introduce noise and redundancy if concepts are too similar or irrelevant.
Table 14 lists the textual concepts used for the ablation in Figure 14. The sets differ because we prompted the LLM five times to generate concepts for the CBM using the following prompt:
Create unique concepts (>100) for a concept-bottleneck model for
the dataset ’Breakfast Actions’. Return them in this format:
"prepare coffee, grind beans, ..."
Both the number and the variety of concepts vary across calls. Despite this variability, model performance remains stable, demonstrating MoTIF’s robustness to different concept sets. The same prompting procedure was repeated for the other datasets (e.g. HMDB51, UCF101, Something-Something V2).
Table 13: Concept-Set Ablation on SSv2. Top-1 accuracy (%) with PE-L/14 evaluated on the four additional concept sets.
Concept Set MoTIF MoTIF-ST Global CBM
(1) Nouns only 18.0 20.3 14.8
(2) Verbs only 24.4 28.6 21.2
(3) Nouns + Verbs 22.1 24.5 18.2
(4) All concepts 29.8 36.0 26.7
(5) Original set 30.0 35.1 25.5
The ablation yields four observations:
  • Across all concept sets, MoTIF improves over Global CBM, showing that MoTIF’s temporal modeling reliably strengthens concept grounding regardless of the chosen vocabulary.
  • Verb-only concepts (set 2) achieve better performance than the noun-only (set 1) counterpart, indicating that MoTIF is particularly effective at grounding action dynamics that require temporal structure.
  • The curated concept set (set 4) yields the highest performance (36.0% - with MoTIF-ST), suggesting that a balanced vocabulary provides the best trade-off between coverage and specificity.
Table 14: MoTIF Breakfast Concepts. The textual concepts utilized for the ablation of concept influence.
Set Count Concepts
Breakfast Original 223 add, adjust, apple, arrange, assemble, avocado, bacon, bagel, bake, balance, banana, batter, beat, bin, blend, blender, blow, boil, bottle, bowl, bread, brew, brush, burner, butter, button, carry, carton, catch, cereal, chair, cheese, chop, cinnamon, clap, close, coffee, colander, comb, container, cook, cookbook, cool, core, counter, cover, crack, croissant, cucumber, cup, cupboard, cut, cuttingboard, detergent, dish, drag, drain, drizzle, drop, dry, egg, faucet, fill, flame, flip, fold, fork, freezer, fridge, froth, frown, fruit, fry, garbage, gesture, granola, grate, grater, grill, grind, ham, handle, heat, herb, hide, honey, hood, ice, ingredient, insert, jar, juice, kettle, knife, knob, knock, ladle, laugh, leftover, lid, mash, measure, measuringcup, measuringspoon, milk, mix, mug, napkin, nod, onion, open, orange, oven, ovenmitt, pack, package, pan, pantry, pastry, peel, peeler, pick, pinch, pit, place, plate, plug, point, poke, pour, preheat, press, pull, push, put, reach, recipe, recycle, release, remove, reveal, rinse, roll, rotate, sausage, scale, scoop, scramble, scrub, seal, serve, serving, set, shake, shave, sieve, sink, sip, sit, slice, slide, smile, snap, soap, socket, sort, spatula, spin, sponge, spoon, spread, sprinkle, squeeze, stack, stand, start, steam, steep, stir, stirrer, stool, stop, stove, strawberry, sugar, switch, syrup, table, take, tamp, tap, taste, tea, thermometer, throw, tie, tilt, timer, toast, tomato, tongs, toss, towel, tray, turn, twist, uncover, unfold, unscrew, unstack, untie, unwrap, warm, wash, waste, water, wave, whisk, wipe, wring, yogurt, zest, zip
Breakfast Set 2 140 adjust heat, arrange cutlery, bake bread, bake pastry, beat eggs, bend down, bite food, break chocolate, break egg shell, breakfast clock, bubbling liquid, butter toast, chair at table, chew food, chop onion, clean spoon, close carton, close cupboard, close drawer, close fridge, close jar, close microwave, close oven, close tap, cut banana, cut dough, cut sandwich, dice tomato, drip water, drizzle dressing, drizzle honey, drizzle oil, dry dish, dry hands, empty sink, family sitting, flip bread, flip toast, fold mixture, fold napkin, grab fork, grab knife, grab spoon, grate chocolate, grill sandwich, gulp drink, hold bowl, hold glass, hold plate, hold straw, knead dough, lean forward, lick spoon, mash egg, mash potato, melt chocolate, mix salad, mop spill, morning light, move chair, one person eating, open carton, open cupboard, open drawer, open fridge, open microwave, open oven, open tap, peel apple, peel banana, peel egg, peel orange, person standing at stove, place utensil, pour batter, pour carton, preheat oven, press button, put down bowl, put down glass, put down plate, reach cupboard, reach shelf, recycle carton, recycle glass, recycle paper, recycle plastic, rest dough, rinse cup, roast vegetable, roll dough, run water, separate yolk, set napkin, shape dough, shred lettuce, sip drink, sip straw, sit down, sizzling pan, slice cucumber, soap hands, spread batter, spreading butter, spreading jam, spreading topping, sprinkle cheese, sprinkle herbs, sprinkle spices, squeeze lemon, stack plates, stand up from chair, steam rising, stir chocolate, swallow food, take carton, take jar, take package, take utensil, tear package, throw trash, toast bun, toss salad, towel hands, turn knob, turn off kettle, turn off stove, turn on kettle, turn on stove, two people cooking, unwrap sandwich, wash dish, wash hands, water boiling, whisk whites, wipe counter, wipe knife, wipe plate, wrap sandwich, zest lemon
Breakfast Set 3 128 add cinnamon, add fruit topping, add granola, add honey, add ice to blender, add milk, add milk to cereal, add sugar, arrange cutlery, assemble sandwich, bake pastry, beat eggs, blend smoothie, boil potato, boil water, brew coffee, butter toast, check timer, chop herbs, chop onion, chop vegetables, clear table, close cupboard, close fridge, close jar, close oven, cook bacon, cook pancake, cook sausage, core apple, crack egg, cut sandwich, dice vegetables, drain bacon, drizzle honey, drizzle syrup, dry dishes, fill kettle, flip bacon, flip omelette, flip pancake, follow recipe, froth milk, fry egg, grate cheese, grill sandwich, grind coffee beans, heat pan, insert coffee pod, make omelette, mash avocado, mash potato, measure ingredients, mix batter, open cupboard, open egg carton, open fridge, open jar, open oven, operate espresso machine, pack leftovers, peel banana, peel potato, pick up spoon, pit avocado, place cup, place plate, pour batter, pour cereal into bowl, pour coffee into cup, pour hot water, pour milk, pour smoothie, pour syrup, pour yogurt into bowl, preheat oven, prepare coffee, pull espresso shot, put ingredient in fridge, put leftovers in fridge, read recipe, rinse fruit, scramble eggs, serve pancakes, set table, set timer, sip beverage, slice apple, slice avocado, slice bagel, slice banana, slice bread, slice cheese, slice cucumber, slice ham, slice orange, slice strawberries, slice tomato, spread butter, spread cream cheese, spread jam, spread peanut butter, sprinkle sugar, squeeze lemon, steam milk, steep tea, stir beverage, strain smoothie, take cup, take ingredient from fridge, take leftovers out fridge, take plate, toast bagel, toast bread, unscrew lid, unwrap bread, use fork, use knife, use measuring cup, use measuring spoon, use spatula, use tongs, use whisk, warm croissant, wash dishes, wash fruit, whisk eggs, wipe counter
Breakfast Set 4 175 add cereal, add cinnamon, add cocoa powder, add honey, add ice to blender, add milk, add milk to coffee, add pepper, add salt, add sugar to coffee, add sugar to tea, add toppings, adjust seasoning, arrange cutlery, assemble sandwich, beat eggs, blend smoothie, blow on hot food, boil potato, boil water, brew coffee, butter toast, carry plate to table, check food temperature, check timer, chop herbs, chop onion, chop tomato, clean blender, clean counter, clear table, close cupboard, close egg carton, close fridge, close jar, close microwave, close milk carton, close oven, cook bacon, cook sausage, crack egg, cut lemon, cut sandwich, dice vegetables, drain bacon, drizzle honey, drizzle syrup, dry dishes, dry hands, fill kettle, fill pot with water, flip bacon, flip omelette, flip pancake, follow recipe, froth milk, fry egg, grate cheese, grind coffee beans, hold pan lid, insert bread into toaster, insert coffee pod, make omelette, mash potato, measure ingredients, mix batter, open cupboard, open egg carton, open fridge, open jar, open microwave, open milk carton, open oven, open package, operate blender, operate espresso machine, pack leftovers, peel banana, peel orange, peel potato, pick up cup, pick up knife, pit avocado, place pan off stove, place pan on stove, place plate, pour batter, pour cereal into bowl, pour coffee, pour eggs into pan, pour from carton, pour hot water, pour milk into bowl, pour pancake batter, pour smoothie, pour syrup, pour tea, pour yogurt into bowl, preheat oven, prepare coffee, prepare tea, press coffee, pull espresso shot, put leftovers in fridge, reach for ingredient, read recipe, remove bread from toaster, remove lid from pot, rinse fruit, scoop butter, scramble eggs, search for ingredient, season food, serve omelette, serve pancakes, set table, set timer, sip beverage, sit down, slice apple, slice avocado, slice bagel, slice banana, slice bread, slice cheese, slice cucumber, slice fruit, slice ham, slice kiwi, slice orange, slice pancake stack, slice strawberries, slice tomato, spoon yogurt, spread butter, spread cream cheese, spread jam, spread peanut butter, sprinkle granola, sprinkle sugar, squeeze lemon, stand up, start microwave, steam milk, steep tea, stir coffee, stop microwave, strain smoothie, take leftovers out fridge, take plate, taste food, toast bagel, toast bread, turn off kettle, turn off stove, turn on kettle, turn on stove, unscrew lid, unwrap bread, use fork, use french press, use knife, use measuring cup, use oven mitts, use spatula, use spoon, use toaster, use tongs, use whisk, wash blender, wash dishes, wash fruit, wash hands, whisk eggs, wipe counter
Breakfast Set 5 161 add cinnamon, add fruit topping, add granola, add honey, add ice to blender, add milk, add milk to cereal, add nuts, add sugar, adjust seasoning, arrange cutlery, assemble sandwich, bake pastry, beat eggs, blend smoothie, blow on hot food, boil potato, boil water, brew coffee, butter toast, carry plate to table, check timer, chop herbs, chop onion, chop vegetables, clear table, close cupboard, close fridge, close jar, close milk carton, close oven, cook bacon, cook pancake, cook sausage, core apple, core pineapple, crack egg, cut parsley, cut pineapple, cut sandwich, dice vegetables, drain bacon, drain can, drizzle honey, drizzle syrup, dry dishes, dry hands, fill kettle, flip bacon, flip omelette, flip pancake, follow recipe, froth milk, fry egg, grate cheese, grill sandwich, grind beans, heat pan, insert coffee pod, juice orange, make omelette, mash avocado, mash potato, measure ingredients, mix batter, open cupboard, open egg carton, open fridge, open jar, open milk carton, open oven, open package, open tin, operate espresso machine, pack leftovers, peel banana, peel orange, peel potato, pick up cup, pick up spoon, pit avocado, place cup, place pan off stove, place plate, pour batter, pour cereal, pour coffee into cup, pour from carton, pour hot water, pour milk, pour smoothie, pour syrup, pour yogurt, preheat oven, prepare coffee, pull espresso shot, put ingredient in fridge, put leftovers in fridge, reach for ingredient, read recipe, remove lid from pot, rinse fruit, scramble eggs, seal container, serve pancakes, set table, set timer, sip beverage, sit down, slice apple, slice avocado, slice bagel, slice banana, slice bread, slice cheese, slice cucumber, slice ham, slice kiwi, slice lemon, slice melon, slice orange, slice pear, slice strawberries, slice tomato, spread butter, spread cream cheese, spread jam, spread peanut butter, sprinkle sugar, squeeze lemon, stand up, steam milk, steep tea, stir beverage, strain smoothie, take cup, take ingredient from fridge, take leftovers out fridge, take plate, taste food, toast bagel, toast bread, toast nuts, unscrew lid, unwrap bread, unwrap package, use fork, use knife, use measuring cup, use measuring spoon, use oven mitts, use spatula, use tongs, use whisk, warm croissant, wash dishes, wash fruit, wash hands, whisk eggs, wipe counter, zest lemon
E.4 Ablation of Concept Construction
To evaluate whether MoTIF can effectively ground explicitly temporal concepts, we constructed five concept sets for SSv2, the dataset in our benchmark that relies most heavily on temporal reasoning: (1) nouns only, (2) verbs only, (3) nouns combined with verbs, (4) the curated concept set used in the main paper, and (5) the union of all concepts (see Table 15). All variants were evaluated using MoTIF with and without space-time transformer architecture. Across all settings, MoTIF consistently outperforms Global CBM.
Table 15: MoTIF Breakfast Concepts. Textual concepts used in the concept-set ablation.
Set Count Concepts
Nouns-Only Concepts 60 object, container, box, cup, bowl, plate, spoon, knife, fork, chopstick, pen, pencil, paper, book, phone, remote, laptop, keyboard, mouse, bag, backpack, toy, ball, fruit, apple, orange, banana, grape, vegetable, carrot, cucumber, tomato, bottle, can, lid, cap, key, lock, door, window, wall, floor, table, chair, shelf, hand, finger, arm, face, person, background, surface, inside, outside, top, bottom, left, right, upward, downward
Verbs-Only Concepts 135 push, pull, lift, drop, hold, carry, throw, catch, slide, drag, roll, spin, rotate, flip, fold, unfold, wrap, unwrap, tie, untie, fasten, unfasten, tighten, loosen, break, cut, slice, chop, tear, peel, crumple, flatten, bend, stretch, shake, stir, pour, scoop, sprinkle, stack, unstack, assemble, disassemble, open, close, lock, unlock, press, tap, swipe, scroll, zoom in, zoom out, point, touch, wave, clap, knock, snap, swing, juggle, bounce, balance, topple, insert, remove, fill, empty, mix, separate, spill, scatter, gather, cover, uncover, hide, reveal, lean, tilt, climb, crawl, jump, hop, walk, run, sprint, stumble, fall, get up, sit, stand, kneel, crouch, bow, dance, nod, shake head, smile, frown, laugh, cry, shout, whisper, speak, yawn, sneeze, cough, sleep, wake, eat, chew, bite, sip, drink, spit, blow, smell, taste, write, draw, erase, paint, type, click, drag mouse, plug, unplug, connect, disconnect, turn on, turn off, start, stop, accelerate, decelerate
Noun–Verb Concepts 104 lift the box, open the box, close the box, drop the box, push the box, pull the box, carry the cup, pour from the cup, fill the cup, empty the cup, hold the cup, rotate the cup, open the bottle, close the bottle, pour from the bottle, lift the bottle, drink from the bottle, cut the paper, fold the paper, tear the paper, crumple the paper, write on the paper, open the book, close the book, flip the book, read the book, drop the book, type on the keyboard, plug in the laptop, unplug the laptop, open the laptop, close the laptop, click the mouse, drag the mouse, press the key, turn on the phone, turn off the phone, swipe the phone, tap the phone, scroll on the phone, charge the phone, unlock the phone, lock the door, unlock the door, open the door, close the door, lift the chair, move the chair, sit on the chair, stand from the chair, throw the ball, catch the ball, bounce the ball, roll the ball, peel the banana, eat the banana, cut the apple, eat the apple, slice the cucumber, pour the water, stir the soup, mix the ingredients, chop the vegetables, open the can, close the lid, place the lid, remove the lid, stack the boxes, unstack the boxes, wrap the gift, unwrap the gift, tie the rope, untie the rope, press the button, flip the switch, turn the knob, insert the plug, remove the plug, connect the cable, disconnect the cable, shake the bottle, squeeze the bottle, pour the juice, stir the drink, draw on the paper, paint the wall, erase the drawing, fold the towel, open the backpack, close the backpack, pick up the bag, drop the bag, open the window, close the window, wipe the table, clean the floor, open the box with one hand, lift the object, move the object, drop the object, hold the object, place the object, throw the object, pick up the object
All Concepts 391 push, pull, lift, drop, hold, carry, throw, catch, slide, drag, roll, spin, rotate, flip, fold, unfold, wrap, unwrap, tie, untie, fasten, unfasten, tighten, loosen, break, cut, slice, chop, tear, peel, crumple, flatten, bend, stretch, shake, stir, pour, scoop, sprinkle, stack, unstack, assemble, disassemble, open, close, lock, unlock, press, tap, swipe, scroll, zoom in, zoom out, point, touch, wave, clap, knock, snap, swing, juggle, bounce, balance, topple, insert, remove, fill, empty, mix, separate, spill, scatter, gather, cover, uncover, hide, reveal, lean, tilt, climb, crawl, jump, hop, walk, run, sprint, stumble, fall, get up, sit, stand, kneel, crouch, bow, dance, spin dance, nod, shake head, smile, frown, laugh, cry, shout, whisper, speak, yawn, sneeze, cough, sleep, wake, eat, chew, bite, sip, drink, spit, blow, smell, taste, write, draw, erase, paint, type, click, drag mouse, plug, unplug, connect, disconnect, turn on, turn off, start, stop, accelerate, decelerate, pretend to push, pretend to pull, pretend to pour, pretend to eat, pretend to drink, pretend to throw, pretend to catch, pretend to type, pretend to swipe, pretend to scroll, pretend to climb, pretend to fall, pretend to hug, pretend to kiss, pretend to wave, pretend to play guitar, pretend to drive, pretend to steer, pretend to read, pretend to sleep, pretend to wake, pretend to write, pretend to draw, pretend to paint, pretend to clean, pretend to cook, pretend to stir, pretend to measure, pretend to weigh, pretend to look around, pretend to search, pretend to point, pretend to balance, pretend to open, pretend to close, pretend to lock, pretend to unlock, pretend to kick, pretend to punch, pretend to block, pretend to dodge, pretend to jump rope, pretend to row, pretend to paddle, pretend to shoot arrow, pretend to load gun, pretend to fire gun, pretend to throw ball, pretend to dribble, pretend to shoot basket, pretend to swing bat, pretend to serve, pretend to catch fish, pretend to steer wheel, pretend to honk, pretend to use controller, pretend to play piano, pretend to play drums, pretend to dance, pretend to sing, pretend to clap, pretend to salute, pretend to bow, pretend to shake hands, pretend to hug, pretend to kiss, lift the box, open the box, close the box, drop the box, push the box, pull the box, carry the cup, pour from the cup, fill the cup, empty the cup, hold the cup, rotate the cup, open the bottle, close the bottle, pour from the bottle, lift the bottle, drink from the bottle, cut the paper, fold the paper, tear the paper, crumple the paper, write on the paper, open the book, close the book, flip the book, read the book, drop the book, type on the keyboard, plug in the laptop, unplug the laptop, open the laptop, close the laptop, click the mouse, drag the mouse, press the key, turn on the phone, turn off the phone, swipe the phone, tap the phone, scroll on the phone, charge the phone, unlock the phone, lock the door, unlock the door, open the door, close the door, lift the chair, move the chair, sit on the chair, stand from the chair, throw the ball, catch the ball, bounce the ball, roll the ball, peel the banana, eat the banana, cut the apple, eat the apple, slice the cucumber, pour the water, stir the soup, mix the ingredients, chop the vegetables, open the can, close the lid, place the lid, remove the lid, stack the boxes, unstack the boxes, wrap the gift, unwrap the gift, tie the rope, untie the rope, press the button, flip the switch, turn the knob, insert the plug, remove the plug, connect the cable, disconnect the cable, shake the bottle, squeeze the bottle, pour the juice, stir the drink, draw on the paper, paint the wall, erase the drawing, fold the towel, open the backpack, close the backpack, pick up the bag, drop the bag, open the window, close the window, wipe the table, clean the floor, open the box with one hand, lift the object, move the object, drop the object, hold the object, place the object, throw the object, pick up the object, object, container, box, cup, bowl, plate, spoon, knife, fork, chopstick, pen, pencil, paper, book, phone, remote, laptop, keyboard, mouse, bag, backpack, toy, ball, fruit, apple, orange, banana, grape, vegetable, carrot, cucumber, tomato, bottle, can, lid, cap, key, lock, door, window, wall, floor, table, chair, shelf, hand, finger, arm, face, person, other, background, surface, inside, outside, top, bottom, left, right, upward, downward, hot, cold, wet, dry, clean, dirty, empty, full, broken, fixed, smooth, rough, heavy, light, fragile, durable, rollable, stackable, squeezable, pourable, spillable, openable, closeable, edible, drinkable